### v1

Abstract
The integration of Large Language Models (LLMs) with Knowledge Representation Learning (KRL) signifies a pivotal advancement in the field of artificial intelligence, enhancing the ability to capture and utilize complex knowledge structures. This synergy leverages the advanced linguistic and contextual understanding capabilities of LLMs to improve the accuracy, adaptability, and efficacy of KRL, thereby expanding its applications and potential. Despite the increasing volume of research focused on embedding LLMs within the domain of knowledge representation, a thorough review that examines the fundamental components and processes of these enhanced models is conspicuously absent. Our survey addresses this by categorizing these models based on three distinct Transformer architectures, and by analyzing experimental data from various KRL downstream tasks to evaluate the strengths and weaknesses of each approach. Finally, we identify and explore potential future research directions in this emerging yet underexplored domain, proposing pathways for continued progress.
keywords:
Knowledge graph, Large language model, Representation learning
1 Introduction
Large language models (LLMs) (e.g., BERT 18 , LLaMA 59 ) which represents a direction of ever-increasing model sizes pre-trained on larger corpora, have demonstrated powerful capabilities in solving natural language processing (NLP) tasks, including question answering 99 , text generation 100 and document understanding 101 . There are no clear and static thresholds regarding the model sizes. Early LLMs (e.g., BERT, RoBERTa) adopt an encoder architecture and show capabilities in text representation learning and natural language understanding. In recent years, more focus has been given to larger encoder-decoder 102 or decoder-only 103 architectures. As the model size scales up, such LLMs have also shown reasoning ability and even more advanced emergent ability 104 , exposing a strong potential for Artificial General Intelligence (AGI).
This inflection point, with the arrival of LLMs, marks a paradigm shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge. As a popular approach for explicit knowledge representation, KGs are now widely investigated for the combination with Transformer-based LLMs, including pre-trained masked language models (PLMs) like BERT and RoBERTa, and more recent generative LLMs like the GPT series and LLaMA. Some works use LLMs to augment knowledge graph representation learning. In this survey, considering three directions, encoder-based methods, encoder-decoder-based method, and decoder-based methods. We present a better understanding of the shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge.
Cao et al. 22 and Biswas et al. 40 discuss recent advancements in knowledge graph representation learning, yet they inadequately address aspects related to the integration with large models. Pan et al. 42 and Pan et al. 43 explore the combination of knowledge graphs and large models, specifically addressing LLM4KG and KG4LLM; however, they provide limited coverage of representation learning. Consequently, there is currently no comprehensive review article dedicated to outlining the latest developments in the field of knowledge graph representation learning.
Contributions.   The notable contributions of this survey are summarized as follows:
  • Categorization of KRL. We systematically summarize the knowledge representation learning where large language models can be adopted into: encoder-based, encoder-decoder-based, and decoder-based methods.
  • Systematic Review of Techniques. We provide the most comprehensive overview of large language models on knowledge graph representation learning techniques. For different methods, we summarize the representative models, provide detailed illustrations, and make necessary comparisons.
  • Future Directions. We delve into the foundational principles of language models on graphs and propose six prospective avenues for future exploration.
Organization.   The structure of this survey is organized as follows. Section 2 introduce the foundation by introducing the foundational concepts pertinent to this review, including Knowledge Graphs, Large Language Models, and Text-based Representation Learning. Sections 3 investigates the core methodologies employed in current research, categorizing them into three types of Transformer, respectively. Each of these sections further explores various sub-methodologies. In Section 4, we discuss the advantages of each type of approach from the performance evaluation results of each downstream task. Section 5 outlines potential future research directions that align with the discussed categories, proposing areas that promise significant advancements. The conclusion in Section 6 synthesizes the insights drawn from the review and highlights the implications for future research.
2 Preliminaries
This section provides formal definitions and relevant notational conventions (as shown in Tab. LABEL:notation_table) used in this survey.
2.1 Knowledge Graph
A KG $\mathcal{G}$ is a labelled directed graph, which can be viewed as a set of knowledge triples $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times(\mathcal{E}\cup\mathcal{L})$, where $\mathcal{E}$ is the set of nodes, corresponding to entities (or resources), $\mathcal{R}$ is the set of relation types (or properties) of the entities, and $\mathcal{L}$ is the set of literals. An entity represents a real-world object or an abstract concept. Often the labels of entities and relations are chosen to be URIs or IRIs (Internationalised Resource Identifiers).
Given a KG $\mathcal{G}$, we call $(e_{h},r,e_{t})\in\mathcal{T}$ a triple, where $e_{h}\in\mathcal{E}$ is the subject, $r\in\mathcal{R}$ is the relation, and $e_{t}\in\mathcal{E}\cup\mathcal{L}$ is the object. The subject is also called the head entity, and an object $e_{t}\in\mathcal{E}$ may be referred to as the tail entity. Triples with literals as objects, i.e., $e_{t}\in\mathcal{L}$ are known as attributive triples. In this survey, we use the notation $\langle e_{h},r,e_{t}\rangle$, with angle brackets, to indicate a triple.
Depending on the nature of the objects in a triple, one may distinguish two main kinds of relations:
[…]
Text literals store information in the form of free natural language text and are often used for labels, entity descriptions, comments, titles, and so on.
2.2 Large Language Model
Large language models (LLM) are built upon the pretraining technique, which aims to train a general model using large amounts of data and tasks that can be fine-tuned easily in different downstream applications. The concept of pretraining originates from transfer learning in computer vision (CV) tasks. Recognizing the effectiveness of pretraining in CV, researchers have extended pretraining techniques to other domains, including natural language processing (NLP).
When applied to NLP, well-trained language models capture rich knowledge beneficial for downstream tasks, such as understanding long-term dependencies, hierarchical relationships, and more. One of the significant advantages of pretraining in NLP is that training data can be derived from any unlabeled text corpus, providing an essentially unlimited amount of training data.
Early pretraining methods in NLP, such as the Neural Language Model (NLM) and Word2vec, were static, meaning they could not adapt effectively to different semantic environments. Static models learn fixed word representations, which limits their ability to handle context variations. To overcome this limitation, dynamic pretraining techniques were developed, leading to Pretrained Language Model (PLM) like BERT (Bidirectional Encoder Representations from Transformers) and XLNet. These models dynamically adjust word representations based on context, significantly enhancing their adaptability and performance across various tasks.
BERT models the conditional probability of a word given its bidirectional context, also named masked language modeling (MLM) objective:
$\mathbb{E}_{\mathcal{S}\sim\mathcal{D}}\left[\sum_{s_{i}\in\mathcal{S}}\log p(s_{i}\mid s_{1},\ldots,s_{i-1},s_{i+1},\ldots,s_{N_{\mathcal{S}}})\right]$ (1)
where $\mathcal{S}$ is a sentence sampled from the corpus $\mathcal{D}$, $s_{i}$ is the $i$-th word in the sentence, and $N_{\mathcal{S}}$ is the length of the sentence. BERT utilizes the Transformer architecture with attention mechanisms as the core building block. In the vanilla Transformer, the attention mechanism is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (2)
where $Q,K,V\in\mathds{R}^{N_{S}\times d_{k}}$ are the query, key, and value vectors for each word in the sentence, respectively. Following BERT, other masked language models are proposed, such as RoBERTa, ALBERT, and ELECTRA, with similar architectures and objectives of text representation.
Although the original Transformer paper was experimented on machine translation, it was not until the release of GPT-2 that language generation (aka. causal language modeling) became impactful on downstream tasks. Causal language modeling is the task of predicting the next word given the previous words in a sentence. The objective of causal language modeling is defined as:
$\mathbb{E}_{\mathcal{S}\sim\mathcal{D}}\left[\sum_{s_{i}\in\mathcal{S}}\log p(s_{i}\mid s_{1},\ldots,s_{i-1})\right]$ (3)
A prime example of the application of LLM in NLP is ChatGPT, which is fine-tuned from the generative pretrained transformer GPT-3.5. GPT-3.5 was trained on a diverse blend of text and code, making it adept at a wide range of tasks from conversation to code generation. ChatGPT leverages reinforcement learning from human feedback (RLHF), which has emerged as a promising approach to align large language models with human intent. This alignment is achieved through iterative tuning based on human feedback, ensuring the responses are more relevant and accurate.
Figure 1: The history and evolution of LLM.
We would like to point out that the word “large” in LLM is not associated with a clear and static threshold to divide language models. “Large” actually refers to a direction in which language models are inevitably evolving, and larger foundational models tend to possess significantly more representation and generalization power. Hence, we define LLMs to encompass both medium-scale PLMs, such as BERT, and large-scale LMs, like GPT-4, as suggested.
2.3 Knowledge Representation Learning
The models for knowledge representation learning can be categorized into three main types: Linear Models, Convolutional Neural Networks (CNNs), and Recurrent Neural Networks (RNNs).
Simple linear models can be used to generate text-based representations. Description-embodied knowledge representation learning (DKRL) and Joint(BOW) use continuous bag-of-words (CBOW) to encode keywords extracted for entity descriptions. The vectors of these keywords are summed up to form text-based entity representations. However, CBOW has limitations as it treats all keywords equally and neglects word order in the description. Veira et al. leverage Wikipedia entries for entity descriptions and generate relation-specific weighted word vectors (WWV) for entities. WWV assigns different importance to words in the description based on frequency and their relationship with the relation of the word. Matrix A records the number of occurrences of each word in the description, while matrix B captures relevance.
Convolutional neural networks (CNNs) are effective for encoding entity descriptions and textual mentions, learning deep expressive features from textual information. DKRL(CNN) assumes that word order in entity descriptions involves implicit relations between entities that KGs may omit. It uses a five-layer CNN to discover these implicit relations. The vector embeddings of words, excluding stop words, are input to the CNN, with max-pooling and mean-pooling operations used to manage parameter space and filter noise. The non-linear output layer constructs the text-based entity representation.
Recurrent neural networks (RNNs) capture long-term relational dependencies. Sequential Text-embodied Knowledge Representation Learning (STKRL) extracts reference sentences for each entity from a corpus, treating entity representation as a multi-instance learning problem. A position-based RNN/LSTM encoder generates sentence-level representations, assigning importance based on cosine similarity between sentence representations and structure-based embeddings. Entity Descriptions-Guided Embedding (EDGE) uses BiLSTM to encode entity descriptions, handling context and word sequences effectively. Pre-trained word embeddings from word2vec are input to BiLSTM, refined iteratively with structure-based embeddings to enhance representation quality.
Table 1: Summary of the embeddings, extended models and combination mechanism in text-based KG embedding.
Models Structual embedding Textual embedding Extended model Combination mechanism WVV $r_{s}\in\mathbb{R}^{d}$ $h^{(r_{s})}_{d},t^{(r_{s})}_{d}\in\mathbb{R}^{d}$ SE /TransE/ $f_{r}(h,t)=-\|h^{(r_{s})}_{d}-r_{s}+t^{(r_{s})}_{d}\|$ DKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE RTKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $r_{d}\in\mathbb{R}^{d}$ TransE $f_{r}(h,t)=f_{s}+f_{d}$ RDRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE STKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE Joint $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE $h=g_{h}\otimes h_{s}+(1-g_{h})\otimes h_{d}$ $t=g_{t}\otimes t_{s}+(1-g_{t})\otimes t_{d}$ $r=r_{s}$ EDGE $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
Table 2: Notations of Concepts.
[…]
Table 3: An overview of various LLM enhanced KRL models.
Model Type Base Model Source Code 2019 ERNIE Encoder BERT https://github.com/thunlp/ERNIE KnowBert Encoder BERT https://github.com/allenai/kb KG-BERT Encoder BERT https://github.com/yao8839836/kg-bert 2020 MTL-KGC Encoder BERT https://github.com/bosung/MTL-KGC Pretrain-KGE Encoder BERT - K-BERT Encoder BERT https://github.com/autoliuweijie/K-BERT 2021 StAR Encoder BERT, RoBERTa https://github.com/wangbo9719/StAR_KGC KG-GPT2 Decoder GPT2 - MEM-KGC Encoder BERT - LaSS Encoder BERT, RoBERTa https://github.com/jhshen95/LASS KEPLER Encoder RoBERTa https://github.com/THU-KEG/KEPLER BLP Encoder BERT https://github.com/dfdazac/blp SimKGC Encoder BERT https://github.com/intfloat/SimKGC 2022 LP-BERT Encoder BERT, RoBERTa - PKGC Encoder BERT, RoBERTa, LUKE https://github.com/THU-KEG/PKGC Nayyeri et al. Encoder BERT https://github.com/ZihaoWang/Hypercomplex-KG-Embedding CoDEx Encoder Sentence-BERT - KGT5 Encoder-Decoder T5 https://github.com/apoorvumang/kgt5 OpenWorld KGC Encoder BERT - Lovelace et al. Encoder BERT, PubMedBERT https://github.com/justinlovelace/LM-KG-Completion kNN-KRL Encoder BERT https://github.com/zjunlp/KNN-KG LMKE Encoder BERT https://github.com/Neph0s/LMKE GenKGC Encoder-Decoder BART https://github.com/zjunlp/PromptKG/tree/main/research/GenKGC KG-S2S Encoder-Decoder T5 https://github.com/chenchens190009/KG-S2S 2023 LambdaKG Encoder-Decoder BERT, BART, T5 https://github.com/zjunlp/PromptKG/tree/main/lambdaKG CSPromp-KG Encoder BERT https://github.com/chenchens190009/CSProm-KG ReSKGC Encoder-Decoder T5 - StructGPT Decoder GPT 3.5 https://github.com/RUCAIBox/StructGPT ReasoningLM Encoder RoBERTa https://github.com/RUCAIBox/ReasoningLM KG-GPT Decoder GPT 3.5 https://github.com/jiho283/KG-GPT Keqing Decoder GPT 3.5, LLaMA 2 - AutoKG Decoder GPT 3.5 https://github.com/zjunlp/AutoKG KG-LLM Decoder ChatGLM, LLaMA 2 https://github.com/yao8839836/kg-llm KoPA Decoder Alpaca, LLaMA, GPT 3.5 https://github.com/zjukg/KoPA 2024 CD Decoder PaLM2 https://github.com/DavidLi0406/Contextulization-Distillation CP-KGC Decoder Qwen, LLaMA 2, GPT 4 https://github.com/sjlmg/CP-KGC KICGPT Decoder GPT 3.5 - KnowledgeNavigator Decoder GPT 3.5, LLaMA 2 -
Figure 2: A taxonomy of LLM enhanced KRL.
3 LLM Enhanced KRL Methods
In this section, we delve into the advanced methodologies for enhancing Knowledge Representation Learning using large language models. The section is structured into three main sections, each focusing on a different category of LLM-based methods. Section 3.1 explores Encoder-based methods, which utilize powerful encoders to generate dense vector representations of knowledge entities. Section 3.2 discusses Encoder-Decoder-based methods that leverage both encoding and decoding processes to capture and generate complex relations within knowledge graphs. Finally, Section 3.3 examines Decoder-based methods, emphasizing the generation and prediction capabilities of decoders to infer and expand knowledge representations.
Figure 3: A flowchart of model evolution.
3.1 Encoder-Based Methods
These works employ the encoder of Transformer to transform knowledge graph elements into textual sequences that are then processed to evaluate triple plausibility or to complete knowledge graphs. The contextual capabilities are central to assessing relations and entity attributes in the graph structure.
3.1.1 Multi-Level Encoder
The typical approach is illustrated in Fig. 4. This category primarily focuses on employing BERT or similar encoder-based pre-trained language models to convert knowledge graph elements into textual sequences, which are then used to assess the plausibility of triples. The common method is to use BERT deep contextual embeddings to enhance the representation of entities and relations, thereby improving accuracy in predicting plausible triples.
Figure 4: An overview of multi-level encoder method.
$\text{Input}=[CLS]+\text{Tokenize}(h)+[SEP]+\text{Tokenize}(r)+[SEP]+\text{Tokenize}(t)+[SEP]$ (4)
where [CLS] and [SEP] are special tokens used by BERT, and $\text{Tokenize}(\cdot)$ represents the tokenization process.
The classification task is to determine if the concatenated sequence correctly represents a valid triple from the KG. A logistic sigmoid function computes the probability of the triple being valid:
$s_{\tau}=\sigma(\mathbf{C}\mathbf{W}^{T})$ (5)
Here, $\mathbf{C}$ is the final hidden state of the [CLS] token, $\mathbf{W}$ is a learnable parameter matrix, and $\sigma$ denotes the sigmoid function.
The training uses a cross-entropy loss function to handle both positive and negative examples of triples:
$L=-\sum_{\tau\in D^{+}\cup D^{-}}\left(y_{\tau}\log(s_{\tau,0})+(1-y_{\tau})\log(s_{\tau,1})\right)$ (6)
where $D^{+}$ and $D^{-}$ represent the sets of positive and negative triples, respectively, and $y_{\tau}$ is the label (0 or 1) of the triple.
The use of BERT allows the model to leverage contextualized text embeddings, improving the quality of predictions compared to traditional methods that might only use structural information of the KG.
Pretrain-KGE 74 enriches traditional Knowledge Graph Embeddings (KRLs) using BERT pre-trained embeddings, aiming to leverage linguistic context for better entity and relation representation. The method is beneficial in contexts where semantic richness is critical, but it relies heavily on the quality and scope of the training data.
StAR 88 introduces a hybrid model combining textual encoding with graph embedding techniques, addressing the limitations of both approaches by modeling graph elements and their spatial relations. It benefits from enhanced representational efficiency but might struggle with complexity and scalability due to the dual nature of the encoding.
MEM-KGC 63 utilizes a Masked Language Model approach for predicting masked entities in a knowledge graph, enhancing embeddings for better performance on KGC tasks. Its strength lies in handling unseen entities during testing, though it might introduce biases from the pre-training data used in the MLM.
The advantage of this approach lies in leveraging the robust linguistic capabilities of BERT, which can capture complex patterns in text data that are beneficial for knowledge graph completion. However, the reliance on textual representation may limit the model’s ability to capture purely structural relationships inherent in knowledge graphs, potentially leading to performance bottlenecks when dealing with entities that have poor textual descriptions.
3.1.2 Score Function Variation
This category focuses on utilizing pre-trained language models across multiple tasks simultaneously, such as link prediction, entity resolution, and relation classification. Techniques often involve fine-tuning a shared language model on KG-specific data, employing strategies that allow the model to leverage learned representations more effectively. The process of the typical method is depicted in Fig. 5.
Figure 5: An overview of score function variation method.
Papers in this subcategory explore the enhancement of knowledge graph embeddings through multi-task learning frameworks and fine-tuning processes. These approaches often integrate additional tasks like relation prediction and relevance ranking with the primary link prediction task, aiming to create a more holistic learning scenario that captures various aspects of knowledge graphs.
Each KG triplet $(h,r,t)$ is represented by natural language descriptions of the head $h$, relation $r$, and tail $t$. These descriptions are tokenized into sequences $T_{h}$, $T_{r}$, and $T_{t}$. A pre-trained language model (LM) encodes these sequences. The output embeddings for the head, relation, and tail are obtained by mean pooling the LM outputs for respective tokens. The semantic embedding process uses the formula:
$h=\text{MeanPool}(\text{LM}(T_{h}))$ (7)
$r=\text{MeanPool}(\text{LM}(T_{r}))$ (8)
$t=\text{MeanPool}(\text{LM}(T_{t}))$ (9)
where $T_{h},T_{r},T_{t}$ are token sequences for the head, relation, and tail, respectively.
After semantic embedding, a structured loss is optimized to align the embeddings with the KG structure. This loss is designed to make the embedding of a true triplet $(h,r,t)$ such that $h+r$ approximates $t$.
The structure loss is defined by the scoring function:
$f(h,r,t)=b-\frac{1}{2}\|h+r-t\|^{2}$ (10)
where $b$ is a margin parameter. The loss aims to minimize the Euclidean distance between the predicted and actual tail entity in the embedding space.
The model estimates the probability of a triplet being correct based on the structured score function. To efficiently handle large KGs, negative sampling is employed, which involves generating negative triplets by corrupting either the head or the tail entity.
MTL-KGC 66 integrates multiple learning tasks (link prediction, relation prediction, relevance ranking) to enhance KGC using a shared BERT layer. Multi-task learning helps in leveraging linguistic information across different tasks, enhancing overall learning efficiency. However, it may suffer from task interference where one tasks learning negatively impacts another.
SimKGC 86 introduces an efficient contrastive learning strategy with advanced negative sampling for text-based KGC. It focuses on hard negatives to enhance model training, which improves learning efficiency but may need extensive tuning to optimize the negative sampling strategy.
KEPLER 47 fuses the capabilities of PLMs with traditional KE techniques, offering dual optimization on both factual and linguistic fronts. This integration allows for richer representations, though the complexity of joint optimization could lead to challenges in model convergence.
The multi-task learning approach helps in capturing a more comprehensive semantic understanding by tackling multiple aspects of knowledge graph completion simultaneously. This can lead to models that generalize better on unseen data. However, these methods can be computationally expensive and complex to tune, as they require balancing multiple learning objectives which might conflict.
3.1.3 Negative Strategy
As shown in Fig. 6, some integrate textual information encoded by LLMs with structured knowledge graph embeddings. This fusion approach enhances the ability to handle both the semantic nuances captured through text and the explicit relational data represented in knowledge graphs.
Figure 6: An overview of negative strategy method.
Research in this area focuses on hybrid models that combine the strengths of both textual and structural data. Methods typically involve using language models for initial encoding followed by integration with graph embedding techniques, which help in capturing both textual nuances and relational structures within the graph.
Each entity $v$ and relation $e$ in the graph $\mathcal{G}$ is embedded into a $4\mathbf{D}$-dimensional hypercomplex space, capturing both the structural and textual information:
$v=\mathbf{s}_{v}+\mathbf{x}_{v_{1}}\mathbf{i}+\mathbf{y}_{v_{2}}\mathbf{j}+\mathbf{z}_{v_{3}}\mathbf{k}$ (11)
$e=\mathbf{s}_{e}+\mathbf{x}_{e_{1}}\mathbf{i}+\mathbf{y}_{e_{2}}\mathbf{j}+\mathbf{z}_{e_{3}}\mathbf{k}$ (12)
where $\mathbf{s}_{v}$ and $\mathbf{s}_{e}$ represent the structural embeddings, and $\mathbf{x}_{v_{1}},\mathbf{y}_{v_{2}},\mathbf{z}_{v_{3}}$, etc., are embeddings derived from textual information.
The compatibility of a triplet $(v,e,w)$ in $\mathcal{G}$ is scored using a function that measures the alignment of the Dihedron product of the embeddings:
$f(v,e,w)=\mathbf{Re}\left((\overline{v\otimes e})\otimes w\right)$ (13)
where $\otimes$ denotes the Dihedron product, and $\overline{v\otimes e}$ represents the conjugate of the product.
The method stands out by efficiently integrating diverse information sources and leveraging the mathematical properties of hypercomplex spaces to enhance the semantic richness of knowledge graph embeddings.
CoDEx 28 guides traditional KRL models using embeddings derived from LLMs, enhancing the link prediction capabilities. It benefits from the textual richness of LLMs but depends heavily on the initial quality of the language model embeddings.
OpenWorld KGC 72 presents a unified learning framework that combines MLM with multi-task learning to integrate new entities into KGs effectively. It excels in handling unconnected entities but might struggle with data sparsity issues.
Lovelace et al. 61 optimizes PLMs for KGC by decoupling querying and candidate retrieval embeddings, using supervised and unsupervised techniques. It enhances embedding suitability for KGC, but maintaining scalability and efficiency can be challenging.
The integration of textual and structural embeddings allows for a richer representation of knowledge graphs, potentially improving accuracy in tasks such as link prediction and entity resolution. Nevertheless, the complexity of these models can lead to challenges in training stability and increased computational demands.
3.2 Encoder-Decoder-Based Methods
In this category, LLMs are employed as sequence-to-sequence generators in knowledge graph completion. The models receive textual sequences of query triples and directly generate the text of the tail entity, employing the full generative capabilities of language models to predict missing elements of the graph.
3.2.1 Multi-Level Encoder
Recent works explore advanced contextual embedding strategies and contrastive learning methods to refine knowledge graph completion techniques. The use of rich contextual embeddings from language models combined with innovative training techniques like contrastive loss helps in distinguishing between correct and incorrect triples more effectively. Fig. 7 outlines the steps involved in the conventional method.
Figure 7: An overview of multi-level encoder method.
This subgroup applies contextual embeddings and contrastive learning techniques to knowledge graph completion. Contextual methods leverage language models to generate rich, nuanced embeddings, while contrastive approaches focus on distinguishing between correct and incorrect triples through advanced sampling strategies and loss functions.
For discrimination-based methods, the input sequence for a KG triplet $(v,e,w)$ is prepared as:
$\mathbf{X}_{\text{pair}}=[CLS]\,\mathbf{x}_{v}[SEP]\mathbf{x}_{e}[SEP]$ (14)
$\mathbf{X}_{\text{tail}}=[CLS]\,\mathbf{x}_{w}[SEP]$ (15)
where $\mathbf{x}_{v},\mathbf{x}_{e},\mathbf{x}_{w}$ are the text representations of the head entity, relation, and tail entity, respectively.
For generation-based methods, the scoring function for a KG triplet is defined using the output probabilities of the sequence model:
This represents the likelihood of generating the tail entity sequence given the head entity and relation sequences.
This method represents a significant step forward in KRL research by combining the strengths of PLMs with graph representation learning, thereby improving the effectiveness of knowledge extraction and reasoning tasks.
KG-GPT2 49 employs GPT-2 for KGC, treating triples as sentences for sequence classification. It leverages GPT-2’s contextual understanding but may not always capture the structural nuances of KGs effectively.
GenKGC 40 transitions KGC to a Seq2Seq generation problem using PLMs, focusing on efficient entity generation. It reduces inference times but may sacrifice accuracy for speed.
CD 20 distills context-rich information from LLMs into KGC models, enhancing them with detailed text-based contexts. It improves contextual understanding but requires significant computational resources for the distillation process.
Contextual methods benefit from the deep semantic understanding provided by pre-trained language models, which can significantly enhance the interpretability and performance on complex triples. Contrastive methods improve learning efficiency by focusing on hard negatives. However, these techniques can be sensitive to the choice of negatives and require careful tuning of the contrastive loss parameters.
3.2.2 Soft Prompt
This group utilizes sequence-to-sequence and Transformer architectures to perform tasks such as predicting missing entities in triples or completing parts of the knowledge graph. The flexibility of Seq2Seq models is particularly useful for tasks where structured inputs need to be transformed into textual or other structured outputs. The sequence of actions in the typical method is demonstrated in Fig. 8.
Figure 8: An overview of soft prompt method.
Seq2Seq and Transformer-based models in this category utilize sequence-to-sequence frameworks to generate missing entities or relations in a knowledge graph directly. These models are pre-trained on large text corpora and fine-tuned on knowledge graph data, leveraging their ability to generate coherent and contextually appropriate text sequences.
Each entity, relation, and triplet in the KG is treated as a text sequence. This allows for the use of textual prompts in LLMs for prediction tasks.
$\text{Triple}=(v,e,w)$ (16)
[…]
For triple classification task, which involves determining if a given triple $(v,e,w)$ is correct (valid) or not, the prompt example is:
[…]
For triple classification task, Given $v$ and $w$, predict $e$,
[…]
For entity prediction task, depending on the missing entity, prompts are created to predict either $v$ or $w$, tbe prompt for predicting tail entity $w$ is:
[…]
The prompt for predicting head entity $v$ is:
[…]
Finally, uses instruction tuning with prompts and expected responses to fine-tune the models. This aligns the predictions with the factual correctness required for KG tasks.
KGT5 52 is pre-trained and fine-tuned on Knowledge Graph Completion (KGC) and Knowledge Graph Question Answering (KGQA) tasks. The model employs regularization to maintain link prediction accuracy. The advantage of this method is its flexibility in handling both KGC and KGQA, making it a versatile solution. However, the complexity of training and fine-tuning a Seq2Seq model can be computationally expensive and time-consuming.
KG-S2S 51 adopts a Seq2Seq generative framework to manage various types of KGC tasks by translating graph structures into flat text representations. This approach allows for unified processing but might suffer from a loss of structural information, which is critical for accurate graph completion.
CSPromp-KG 32 leverages Conditional Soft Prompts to guide pre-trained language models in balancing the integration of structural and textual information for KGC. The method enhances model performance by ensuring a more nuanced understanding of graph structures. The major drawback is the potential overfitting to specific knowledge graph structures, which may limit the generalizability.
The strength of Seq2Seq models lies in their flexibility and effectiveness in generating textual output that adheres to the structural rules of knowledge graphs. However, these models can be prone to generating implausible links if not properly constrained or if the training data is not representative enough of the test scenarios.
3.3 Decoder-Based Methods
In these studies, LLMs are adapted to enhance reasoning and inference capabilities over knowledge graphs, particularly for complex question answering and fact verification tasks. Techniques include adapting attention mechanisms to focus on relevant subgraphs and employing reasoning strategies that leverage structured graph data.
3.3.1 Multi-Level Encoder
Fig. LABEL:fig2.2 illustrates the workflow of this approach.
This category includes models that enhance the capabilities of LLMs in complex reasoning and question answering tasks over knowledge graphs. Techniques involve structuring the input to LLMs in a way that mimics reasoning or using retrieval-augmented generation to enhance the context available for decision-making.
The approach starts by converting a relevant subgraph $\mathcal{G}_{q}$ from the knowledge graph into a sequence using a breadth-first search (BFS)-based serialization. This serialized form is then processed by the PLM. The serialized subgraph $\mathcal{G}_{q}$ is denoted as:
$\mathcal{SG}_{q}=\{v,e,w,\ldots\}$ (21)
where $v$, $e$, and $w$ are nodes and edges in the subgraph.
As for subgraph-aware self-attention, the mechanism is designed to mimic the processing of a GNN, enabling the model to understand the structure of the subgraph and reason over it. It utilizes attention mechanisms to model the relations and interactions between different entities (nodes) and relations (edges) within the subgraph. The attention process can be represented as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V$ (22)
Here, $Q$, $K$, and $V$ are query, key, and value matrices, respectively, and $M$ is a mask matrix that directs attention flow to maintain structural integrity.
The PLM is adapted to handle the serialized subgraph and question data simultaneously, facilitating deep interaction and knowledge sharing between the question context and the graph structure. This integration enables the PLM to perform reasoning tasks directly on the subgraph structure, harnessing both the rich semantic understanding of the PLM and the structured reasoning capabilities typically associated with GNNs.
The PLM undergoes adaptation tuning using a dataset created from synthesized questions and subgraphs. This adaptation helps the PLM better understand and process the unique format of KG data. The fine-tuning phase optimizes the PLM on downstream question answering tasks over KGs, improving accuracy and efficiency in real-world applications.
ReSKGC 81 introduces a retrieval-enhanced Seq2Seq model that uses a retrieval module to enhance the generation of missing entities by fetching semantically relevant triplets. This method improves accuracy and relevance but relies heavily on the effectiveness of the retrieval process, which can be a bottleneck if not optimized properly.
StructGPT 88 employs an Iterative Reading-then-Reasoning framework to enhance LLMs’ reasoning capabilities over structured data like knowledge graphs. It significantly improves reasoning performance, but the iterative process can be resource-intensive and slow, particularly with complex queries.
These models excel in tasks requiring deep reasoning and can handle nuanced queries effectively by leveraging both the structured information from knowledge graphs and the unstructured textual data processed by LLMs. The main drawback is their potential computational inefficiency and the complexity of integrating structured reasoning with generative language models.
3.3.2 Prompt Engineering
Research in this area focuses on integrating comprehensive contextual information and structured knowledge into the training and operation of LLMs. This enables the models to better understand and generate responses based on both the textual descriptions and the underlying graph structures of the knowledge graph. The conventional method flowchart can be seen in Fig. 9.
Figure 9: An overview of prompt engineering method.
Methods under this heading focus on integrating extensive contextual information and structured knowledge graph data within the processing capabilities of LLMs. This often involves using knowledge prompts or other forms of guided input to enhance the LLM’s understanding of the graph structure and semantics.
Complex questions are decomposed into simpler sub-questions based on predefined templates. This is achieved through a slot-filling mechanism driven by LLMs, formalized as:
$\{\mathcal{S}_{i,t}\}_{t=1}^{T}=\mathrm{LLM}(\mathcal{Q}_{i})$ (23)
where $\mathcal{Q}_{i}$ is the initial complex question, and $\mathcal{S}_{i,t}$ represents the decomposed sub-questions.
For each decomposed sub-question, the framework retrieves relevant entities and facts from the KG using predefined logical chains that correspond to the sub-questions, the retrieval process is summarized as:
$\mathcal{C}_{t}=\bigcup_{l=1}^{L_{i}}\{(s,r,o)_{l}\mid(s,r,o)_{l}\in\mathcal{G}\}$ (24)
where $(s,r,o)_{l}$ are the triples retrieved based on the logical chains, with $s$, $r$, and $o$ denoting the subject, relation, and object, respectively.
The candidate reasoning stage involves selecting the correct answers from the retrieved candidates by evaluating their relevance to the sub-questions, expressed using the following function:
$\mathcal{A}_{i}=\mathrm{LLM}(\mathcal{Q}_{i,t}\mid\mathcal{C}_{t})$ (25)
where $\mathcal{A}_{i}$ are the answers derived for sub-question $\mathcal{Q}_{i,t}$, based on candidates $\mathcal{C}_{t}$.
Finally, the responses from all sub-questions are aggregated to construct a comprehensive answer to the original question, which formalized as:
$\text{Response}=\text{Generate}(\{\mathcal{A}_{i}\}_{i=1}^{N})$ (26)
where the function Generate compiles the individual answers into a final coherent response.
KoPA 56 introduces the Knowledge Prefix Adapter (KoPA), which enhances the integration of structural KG information into LLMs for better structural-aware reasoning. While it improves reasoning accuracy, the adaptation of LLMs to specific knowledge graph structures could limit the model versatility across different KGs.
CP-KGC 31 explores the use of zero-shot large language models to enhance text-based KGC without fine-tuning. The advantage is the minimal training requirement, but the reliance on zero-shot capabilities may result in less precise or contextually inappropriate text generation.
KICGPT 53 integrates a triple-based knowledge graph completion retriever with a large language model to address information scarcity in long-tail entities. This method effectively utilizes in-context learning strategies, but its success heavily depends on the quality and relevance of the retrieved knowledge.
These approaches allow LLMs to leverage their pre-trained knowledge effectively by guiding them with structured inputs, enhancing their performance on knowledge-intensive tasks. However, the success of these methods heavily depends on the quality and the extent of the integration between text and graph data, which can be challenging to optimize.
4 Experiments and Evaluations
In this section, we present a comprehensive suite of experiments and evaluations to assess the performance and effectiveness of the various LLM-enhanced KRL methods discussed previously. We utilize a variety of datasets and metrics to provide a thorough analysis, ensuring that our findings are robust and generalizable. The section is structured to cover the datasets used, the evaluation metrics employed, and the downstream tasks for which the models were tested.
4.1 Datasets
Firstly, we provide detailed information about the datasets utilized in the KRL experiments. These datasets span multiple tasks, including entity typing, relation classification, triple classification, and link prediction, offering a broad evaluation scope. We outline the key statistics for each dataset, such as the number of entities, relations, and instances in the training, development, and test sets, as well as the specific tasks they are designed for. This comprehensive overview sets the stage for understanding the context and challenges addressed in the subsequent experiments.
Table 4: Statistics on Datasets.
Dataset # Entity # Relation # Train # Dev # Test Task Year Source FIGER - - 2,000,000 10,000 563 ET 2015 1 Open Entity - - 2,000 2,000 2,000 ET 2018 2 FewRel 1.0 - 80 8,000 16,000 16,000 RC 2018 3 TACRED - 42 68,124 22,631 15,509 RC 2017 4 WN11 38,696 11 112,581 2,609 10,544 TC 2013 5 WN9 6,555 9 11,741 1,337 1,319 TC 2022 6 FB13 75,043 13 316,232 5,908 23,733 TC 2013 5 FB15K-237N 13,104 93 87,282 7,041 8,226 TC 2022 7 FB15K 14,951 1,345 483,142 50,000 59,071 LP 2013 8 FB15k-237 14,541 237 272,115 17,535 20,466 LP 2015 9 UMLS 135 46 5,216 652 661 LP 2018 10 Nations 14 55 1,592 199 201 LP 2007 11 Diabetes 7,886 67 56,830 1,344 1,936 LP 2022 2 YAGO3-10 103,222 30 490,214 2,295 2,292 LP 2018 10 Wikidata5M (transductive) 4,594,485 822 20,614,279 5,163 5,133 LP 2021 13 Wikidata5M (inductive) 4,594,458 822 20,496,514 6,699 6,894 LP(ZS) CoDEx-S 45,869 68 32,888 1,827 1,828 LP 2020 14 CoDEx-M 11,941 50 185,584 10,310 10,311 LP CoDEx-L 45,869 69 551,193 30,622 30,622 LP NELL-ONE 68,545 822 189,635 1,004 2,158 LP(ZS) 2018 15 WN18RR 40,943 11 86,835 3,034 3,134 LP, LP(ZS) 2018 10
  • FIGER. The FIGER dataset is designed for fine-grained entity recognition. It provides annotations for a wide range of entity types, offering a comprehensive resource for training models to recognize detailed and specific categories beyond the usual named entities like person, organization, and location.
  • Open Entity. Open Entity is a dataset used for entity typing in an open-domain setting. It provides annotations for entities within free text, helping to improve the classification of entities into a broad set of predefined categories, enhancing the ability to understand and process natural language.
  • FewRel. FewRel (Few-shot Relation Extraction) is a benchmark dataset for evaluating the performance of models on relation classification tasks with limited examples. It includes various relation types and supports the development of models that can learn to classify relations from a few instances.
  • TACRED. TACRED (TAC Relation Extraction Dataset) is a large-scale relation extraction dataset created by Stanford. It includes sentences annotated with a wide range of relation types, providing a rich resource for training and evaluating models on the task of extracting relations between entities in text.
  • WN11. WN11 is a subset of the WordNet knowledge graph used for knowledge graph completion tasks. It focuses on 11 relation types and includes a set of entities and their relations, facilitating research in link prediction and knowledge base inference.
  • FB13. FB13 is derived from the Freebase knowledge graph and is used for relation extraction and knowledge base completion tasks. It includes 13 relation types and provides a benchmark for evaluating the ability of models to predict missing links in a knowledge graph.
  • FB15K. FB15K is a widely used benchmark dataset derived from Freebase. It includes a large set of entities and relations, supporting research in knowledge graph completion, entity resolution, and link prediction.
  • FB15k-237. FB15k-237 is a subset of the FB15K dataset, created to address issues of redundancy and test leakage in the original dataset. It excludes inverse relations, making it a more challenging and realistic benchmark for knowledge graph completion tasks.
  • WN18. WN18 is a benchmark dataset based on the WordNet knowledge graph. It contains a subset of entities and relations, often used for evaluating models on the task of link prediction in knowledge graphs.
  • WN9. WN9 is another subset of the WordNet knowledge graph, designed for tasks such as knowledge graph completion and link prediction. It includes a specific set of entities and relations to facilitate focused research in these areas.
  • WN18RR. WN18RR (WN18 Reversed Relations) is a refined version of the WN18 dataset, created to address issues with inverse relations that were present in the original dataset. It is used for benchmarking link prediction and knowledge graph completion models.
  • UMLS. The Unified Medical Language System (UMLS) is a comprehensive dataset used in biomedical and healthcare research. It integrates various health and biomedical vocabularies, providing a rich resource for developing models in medical text mining and knowledge extraction.
  • NELL-One. NELL-One is a few-shot learning dataset derived from the Never-Ending Language Learning (NELL) system. It supports research in few-shot relation learning by providing a challenging benchmark with a variety of relation types annotated from web text.
  • Wikidata5M. Wikidata5M is a large-scale dataset derived from Wikidata, containing millions of entities and their relations. It is used for various tasks, including knowledge graph completion, entity linking, and question answering, providing a comprehensive resource for natural language understanding.
  • Nations. The Nations dataset is a small-scale dataset used for relation prediction and knowledge graph completion. It includes entities representing countries and various relations among them, offering a simplified benchmark for evaluating models.
  • CoDEx. CoDEx (Comprehensive Knowledge Graph Completion and Explanation) is a suite of knowledge graph completion datasets designed to provide diverse and challenging benchmarks. It includes multiple subsets with different relation types and complexities, supporting research in explainable AI and knowledge graph reasoning.
4.2 Metrics
Here, we describe the various metrics used to evaluate the performance of the models on the different tasks. These metrics include accuracy, precision, recall, F1 score, mean rank, mean reciprocal rank, and Hits@K, among others. Each metric provides unique insights into different aspects of model performance, from classification accuracy to ranking effectiveness. By using a diverse set of metrics, we ensure a well-rounded evaluation of the models’ capabilities.
  • Accuracy. Accuracy is a measure of the number of correct predictions made by a model divided by the total number of predictions. It is calculated as the ratio of correctly classified instances to the total instances and is often used as a primary metric for evaluating classification models. Accuracy is useful when the classes are balanced, but it can be misleading when dealing with imbalanced datasets. The accuracy is calculated as:
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (27)
  • Precision. Precision, also known as positive predictive value, measures the proportion of true positive predictions among all positive predictions made by the model. It is calculated as the number of true positives divided by the sum of true positives and false positives. Precision is particularly important in situations where the cost of false positives is high. The precision is calculated as:
    $\text{Precision}=\frac{TP}{TP+FP}$ (28)
  • Recall. Recall, or sensitivity, measures the proportion of true positive predictions among all actual positive instances. It is calculated as the number of true positives divided by the sum of true positives and false negatives. Recall is critical in scenarios where the cost of false negatives is high, such as in medical diagnosis. The recall is calculated as:
    $\text{Recall}=\frac{TP}{TP+FN}$ (29)
  • F1. The F1 score is the harmonic mean of precision and recall, providing a single metric that balances both. It is particularly useful when the dataset is imbalanced, as it considers both false positives and false negatives. The F1 score is calculated as:
    $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (30)
  • Mean Rank. Mean Rank (MR) is a metric used in ranking tasks, such as knowledge graph completion. It represents the average rank position of the correct entity or relation in the predicted list. Lower MR values indicate better performance, as they imply that the correct entities are ranked higher on average. The MR is calculated as:
    $\text{MR}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}_{i}$ (31)
  • Mean Reciprocal Rank. Mean Reciprocal Rank (MRR) is a measure used to evaluate the effectiveness of a ranking algorithm. It is the average of the reciprocal ranks of the correct answers. MRR is particularly useful for assessing the performance of models in information retrieval and question answering tasks. Higher MRR values indicate better performance. The MRR is calculated as:
    $\text{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\text{rank}_{i}}$ (32)
  • Hits@K. Hits@K is a metric that measures the proportion of correct answers that appear in the top K predicted results. It is commonly used in ranking tasks and knowledge graph completion. For instance, Hits@10 means the percentage of correct answers that are within the top 10 predictions. Higher Hits@K values indicate better performance. The Hits@K is calculated as:
[…]
  • Loose Macro. Loose Macro is an evaluation metric used in multi-label classification tasks. It calculates the precision and recall for each label independently and then averages them. This metric gives equal weight to each label, regardless of how many instances of each label are present, providing a balanced view across all labels. The loose macro precision and recall is calculated as:
    Loose Macro $\text{Loose Macro Precision}=\frac{1}{L}\sum_{j=1}^{L}\frac{TP_{j}}{TP_{j}+FP_{j}}$ (34)
    Loose Macro $\text{Loose Macro Recall}=\frac{1}{L}\sum_{j=1}^{L}\frac{TP_{j}}{TP_{j}+FN_{j}}$ (35)
  • Loose Micro. Loose Micro is another metric for multi-label classification, but it aggregates the contributions of all labels to calculate the precision and recall. It treats each instance-label pair equally, providing a more global view of performance by considering the total number of true positives, false positives, and false negatives across all labels. The loose micro precision and recall is calculated as:
    Loose Micro $\text{Loose Micro Precision}=\frac{\sum_{j=1}^{L}TP_{j}}{\sum_{j=1}^{L}(TP_{j}+FP_{j})}$ (36)
    Loose Micro $\text{Loose Micro Recall}=\frac{\sum_{j=1}^{L}TP_{j}}{\sum_{j=1}^{L}(TP_{j}+FN_{j})}$ (37)
4.3 Downstream Tasks
This section covers the specific downstream tasks for which the models were evaluated. These tasks include entity typing, relation classification, relation prediction, triple classification, and link prediction. For each task, we provide a definition, describe its significance, and present the experimental results obtained.
Figure 10: An overview of downstream tasks achievable by each model.
4.3.1 Entity Typing
Definition 1: Entity Typing. ET is the task of assigning predefined type labels to entities within a given text. This process involves identifying the entity mentions in the text and categorizing them into specific types, such as person, organization, location, or more fine-grained types. The goal is to enhance the understanding of entities within the context and improve the performance of downstream tasks like information extraction and knowledge base construction.
Table 5: Experimental statistics under the ET subtask.
Model Year FIGER Open Entity Choi et al. Acc Macro Micro P R F1 P R F1 NFGEC (Attentive) 70 2016 54.53 74.76 71.58 - - - - - - NFGEC (LSTM) 70 2016 55.6 75.15 71.73 68.8 53.3 60.1 - - - UFET 98 2018 - - - 77.4 60.6 68 68.8 53.3 60.1 BERT 18 2018 52.04 75.16 71.63 76.37 70.96 73.56 76.4 71 73.6 RoBERTa 83 2019 - - - 77.4 73.6 75.4 - - - ERNIE 38 2019 57.19 76.51 73.39 78.42 72.9 75.56 78.4 72.9 75.6 KnowBert 55 2019 - - - 78.7 72.7 75.6 78.6 73.7 76.1 KEPLER 47 2021 - - - 77.8 74.6 76.2 - - -
The transformer-based models’ superior performance highlights their effectiveness in ET tasks. However, this comes at the cost of higher computational requirements and complexity. Models like NFGEC, while less powerful, offer faster training and inference times, making them suitable for applications with limited computational resources.
Enhanced pre-training methods (ERNIE, KnowBert) and integrated approaches (KEPLER) offer significant performance improvements by leveraging additional knowledge sources. These models are particularly useful in scenarios where rich contextual understanding is critical. However, their increased complexity and training requirements necessitate careful consideration of resource availability and deployment constraints.
4.3.2 Relation Classification
Definition 2: Relation Classification. RC is the task of determining the semantic relations between a pair of entities within a given text. This involves identifying the entities and classifying the type of relation that exists between them, such as works at, born in, or located in. The purpose is to extract meaningful relational information from text, which can be used in various applications like knowledge graph construction and question answering.
Table 6: Experimental statistics under the RC subtask.
Model Year FewRel 1.0 TACRED P R F1 P R F1 CNN 23 2015 69.51 69.64 69.35 70.3 54.2 61.2 PA-LSTM 73 2017 - - - 65.7 64.5 65.1 C-GCN 21 2018 - - - 69.9 63.3 66.4 BERT_Base 18 2018 85.05 85.11 84.89 67.23 64.81 66 BERT_Large 18 2018 - - - - - 70.1 RoBERTa 83 2019 - - - 70.4 71.1 70.7 MTB 65 2019 - - - - - 71.5 ERNIE 38 2019 88.49 88.44 88.32 69.97 66.08 67.97 KnowBert 55 2019 - - - 71.6 71.4 71.5 KEPLER 47 2021 - - - 71.5 72.5 72
The transformer-based models’ superior performance highlights their effectiveness in RC tasks. However, this comes at the cost of higher computational requirements and complexity. Models like CNN and PA-LSTM, while less powerful, offer faster training and inference times, making them suitable for applications with limited computational resources.
Graph-based approaches like C-GCN provide valuable insights into the syntactic structure, which can be beneficial for specific applications. However, they require high-quality syntactic information and can be less effective in capturing long-range dependencies compared to transformer-based models.
Enhanced pre-training methods (ERNIE, KnowBert) and integrated approaches (KEPLER) offer significant performance improvements by leveraging additional knowledge sources. These models are particularly useful in scenarios where rich contextual understanding is critical. However, their increased complexity and training requirements necessitate careful consideration of resource availability and deployment constraints.
4.3.3 Relation Prediction
Definition 3: Relation Prediction. RP involves inferring the missing relation between two entities in a knowledge graph. Given a pair of entities, the task is to predict the most likely relation that connects them. This task is crucial for completing knowledge graphs and ensuring that they contain comprehensive and accurate relational information.
Table 7: Experimental statistics under the RP subtask.
Model Year FB15K YAGO3-10-100 MR Hits@1 Hits@1 TransE 92 2013 2.5 84.3 - TransR 96 2015 2.1 91.6 - DKRL 36 2016 2 90.8 - TKRL 90 2016 1.7 92.8 - PTransE 76 2015 1.2 93.6 - SSP 87 2017 1.2 - - ProjE 75 2017 1.2 95.7 - KG-BERT 48 2019 1.2 96 - KGT5 52 2022 - - 60 ChatGPT 2023 - - 39 GPT-4 2023 - - 56 LLaMA-7B 59 2023 - - 13 LLaMA-13B 59 2023 - - 1 KG-LLM 50 2023 - - 71
Transformer-based models like KG-BERT and KG-GPT2 leverage the power of pre-trained language models, which can capture intricate semantic relationships and contextual nuances within triples. This results in a higher accuracy of classification compared to traditional methods that primarily rely on structural embeddings.
Models such as KG-LLM demonstrate the ability to adapt to various knowledge graph tasks, benefiting from their extensive pre-training on diverse text corpora. This adaptability allows them to handle different datasets effectively, maintaining high performance across various triple classification benchmarks.
Enhanced transformer models, such as KG-BERT, incorporate external knowledge into their embeddings, further refining their understanding of entity and relation semantics. This integration allows for more accurate predictions and better handling of complex relational patterns.
4.3.4 Triple Classification
Definition 4: Triple Classification. TC is the task of determining whether a given triplet (subject, predicate, object) in a knowledge graph is valid or not. This involves verifying if the relation (predicate) between the subject and object entities holds true. The task helps in maintaining the integrity of knowledge graphs by identifying and filtering out incorrect or unlikely triplets.
Table 8: Experimental statistics under the TC subtask.
Model Year WN11 FB13 UMLS CoDEX-S FB15k-237N Acc Acc Acc Acc Acc NTN 71 2013 86.2 90 - - - TransE 92 2013 75.9 81.5 78.1 72.1 - DistMult 34 2014 87.1 86.2 86.8 66.8 - ComplEx 25 2016 - - 90.8 67.8 65.7 RotatE 84 2019 - - 92.1 75.7 68.5 KG-BERT 48 2019 93.5 90.4 89.7 77.3 56 KG-GPT2 49 2021 85 89 - - - LaSS 57 2021 94.5 91.8 - - - KGT5 52 2022 72.8 66.3 - - - LMKE 60 2022 - 91.7 92.4 - - GPT-3.5 2023 - - 67.6 54.7 60.2 LLaMA-7B 59 2023 21.1 9.1 - - - LLaMA-13B 59 2023 28.1 17.6 - - - KG-LLM 50 2023 95.6 90.2 86 80.3 80.5 KoPA 56 2023 - - 92.6 82.7 77.7
The performance of these models highlights several trade-offs and considerations:
Traditional models like TransE and TransH are simpler and computationally efficient, making them suitable for applications with limited resources. However, their inability to capture complex relationships limits their accuracy on challenging datasets. Transformer-based models (KG-BERT, KG-GPT2) excel in capturing contextual information and complex patterns, achieving high accuracy across diverse datasets. The downside is their significant computational demands and the need for large amounts of training data.
4.3.5 Link Prediction
Definition 5: Link Prediction. LP is the task of predicting missing links between entities in a knowledge graph. Given a partially completed graph, the goal is to infer new links that are likely to exist based on the existing structure and known connections. This task is essential for expanding and enriching knowledge graphs, making them more useful for various applications like recommendation systems and semantic search.
Table 9: Experimental statistics of WN18RR and FB15k-237 datasets under the LP subtask in transductive settings.
Model Year WN18RR FB15k-237 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 2300 24.3 4.3 44.1 53.2 223 27.9 19.8 37.6 47.4 TransH 94 2014 2524 - - - 50.3 255 - - - 48.6 DistMult 34 2014 3704 44.4 41.2 47 50.4 411 28.1 199 30.1 44.6 TransR 96 2015 3166 - - - 50.7 237 - - - 51.1 TransD 91 2015 2768 - - - 50.7 246 - - - 48.4 ComplEx 25 2016 3921 44.9 40.9 46.9 53 508 27.8 19.4 29.7 45 ConvE 29 2018 4464 45.6 41.9 47 53.1 245 31.2 22.5 34.1 49.7 ConvKB 30 2018 2554 24.9 - - 52.5 257 24.3 - - 51.7 R-GCN 82 2018 6700 12.3 8 13.7 20.7 600 16,4 10 18.1 41.7 KBGAN 46 2018 - 21.5 - - 48.1 - 27.7 - - 45.8 RotatE 84 2019 3340 47.6 42.8 49.2 57.1 177 33.8 24.1 37.5 53.3 KBAT 45 2019 1921 41.2 - - 55.4 270 15.7 - - 33.1 CapsE 19 2019 718 41.5 - - 55.9 403 15 - - 35.6 QuatE 78 2019 3472 48.1 43.6 50 56.4 176 31.1 22.1 34.2 49.5 TuckER 97 2019 - 47 44.3 48.2 52.6 - 35.8 26.6 39.4 54.4 HAKE 44 2019 - 49,7 45.2 51.6 58.2 - 34.6 25 38.1 54.2 AttH 17 2020 - 48.6 44.3 49.9 57.3 - 34.8 25.2 38.4 54 REFE 79 2020 - - - - 56.1 - - - - 54.1 GAATs 39 2020 1270 - - - 60.4 187 - - - 65 Complex-DURA 26 2020 57.1 - - - - 56 - - - - DensE 33 2020 3052 49.1 44.3 50.8 57.9 169 34.9 25.6 38.4 53.5 LineaRE 58 2020 1644 49.5 45.3 50.9 57.8 155 35.7 26.4 39.1 54.5 RESCAL-DURA 80 2020 - 49.8 45.5 - 57.7 - 36.8 27.6 - 55 CompGCN 24 2020 - 47.9 44.3 49.4 54.6 - 35.5 26.4 39 53.5 NePTuNe 69 2021 - - - - 55.7 - - - - 54.7 ComplEx-N3-RP 27 2021 - - - - 58 - - - - 56.8 ConE 28 2021 - 49.6 45.3 51.5 57.9 - 34.5 24.7 38.1 54 Rot-Pro 85 2021 - 45.7 39.7 48.2 57.7 - 34.4 24.6 38.3 54 QuatDE 77 2021 1977 48.9 43.8 50.9 58.6 90 36.5 26.8 40 56.3 NBFNet 68 2021 - 55.1 49.7 - 66.6 - 41.5 32.1 - 59.9 KG-BERT 48 2019 97 21.6 4.1 30.2 52.4 153 23.7 16.9 26 42.7 MTL-KGC 66 2020 89 33.1 20.3 38.3 59.7 132 26.7 17.2 29.8 45.8 Pretrain-KGE 74 2020 - 48.8 43.7 50.9 58.6 - 35 25 38.4 55.4 StAR 88 2021 51 40.1 24.3 49.1 70.9 1117 29.6 20.5 32.2 48.2 MEM-KGC 63 2021 - 57.2 48.9 62 72.3 - 34.9 26 38.2 52.4 LaSS 57 2021 35 - - - 78.6 108 - - - 53.3 SimKGC 86 2021 - 66.7 58.8 72.1 80.5 - 33.6 24.9 36.2 51.1 LP-BERT 62 2022 92 48.2 34.3 56.3 75.2 154 31 22.3 33.6 49 Nayyeri et al. 67 2022 - - - - - - 33.6 24.5 37 52.2 KGT5 52 2022 - 54.2 50.7 - 60.7 - 34.3 25.2 - 37.7 OpenWorld KGC 72 2022 - 55.7 47.5 60.4 70.4 - 34.6 25.3 38.1 53.1 Lovelace et al. 61 2022 - 59.1 51.8 61.6 73.5 - 37.5 28.3 40.8 56.4 LMKE 60 2022 79 61.9 52.3 67.1 78.9 141 30.6 21.8 33.1 48.4 GenKGC 40 2022 - - 28.7 40.3 53.5 - - 19.2 35.5 43.9 KG-S2S 51 2022 - 57.4 53.1 59.5 66.1 - 33.6 25.7 37.3 49.8 kNN-KGE 54 2023 - 57.9 52.5 - - - 28 37.3 - - CSPromp-KG 32 2023 - 57.5 52.2 59.6 67.8 - 35.8 26.9 39.3 53.8 GPT-3.5 2023 - - 19 - - - - 23.7 - - LLaMA-7B 59 2023 - - 8.5 - - - - - - - LLaMA-13B 59 2023 - - 10 - - - - - - - KG-LLM 50 2023 - - 25.59 - - - - - - - CD 20 2024 - 57.6 52.6 60.7 67.2 - - - - - CP-KGC 31 2024 - 67.3 59.9 72.1 80.4 - 33.8 25.1 36.5 51.6 KICGPT 53 2024 - 56.4 47.8 61.2 67.7 - 41.2 32.7 44.8 55.4
Table 10: Experimental statistics of UMLS, Wikidata5M, and FB15k-237N datasets under the LP subtask in transductive setting.
Model Year UMLS Wikidata5M FB15k-237N MR Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 1.84 98.9 109370 25.3 17 31.1 39.2 25.5 15.2 30.1 45.9 TransH 94 2014 1.8 99.5 - - - - - - - - - DistMult 34 2014 5.52 84.6 - 25.3 20.9 27.8 33.4 20.9 14.3 23.4 33 TransR 96 2015 1.81 99.4 - - - - - - - - - TransD 91 2015 1.71 99.3 - - - - - - - - - ComplEx 25 2016 2.59 96.7 - 30.8 25.5 - 39.8 24.9 18 27.6 38 DKRL 36 2016 - - 31566 16 12 18.1 22.9 - - - - ConvE 29 2018 1.51 99 - - - - - 27.3 19.2 30.5 42.9 RoBERTa 83 2019 - - 1381597 0.1 0 0.1 0.3 - - - - RotatE 84 2019 - - - 29 23.4 32.2 39 27.9 17.7 32 48.1 QuatE 78 2019 - - - 27.6 22.7 30.1 35.9 - - - - CompGCN 24 2020 - - - - - - - 31.6 23.1 34.9 48 NePTuNe 69 2021 - - - - - - - - - - - ComplEx-N3-RP 27 2021 - 99.8 - - - - - - - - - KG-BERT 48 2019 1.47 99 - - - - - 20.3 13.9 20.1 40.3 MTL-KGC 66 2020 - - - - - - - 24.1 16 28.4 43 StAR 88 2021 1.49 99.1 - - - - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - LaSS 57 2021 1.56 98.9 - - - - - - - - - KEPLER 47 2021 - - 200267 21 17.3 22.4 27.7 - - - - SimKGC 86 2021 - - - 35.8 31.3 37.6 44.1 - - - - LP-BERT 62 2022 1.18 100 - - - - - - - - - KGT5 52 2022 - - - 33.6 28.6 36.2 42.6 - - - - GenKGC 40 2022 - - - - - - - - 18.7 27.3 33.7 KG-S2S 51 2022 - - - - - - - 35.4 28.5 38.8 49.3 CSPromp-KG 32 2023 - - - 38 34.3 39.9 44.6 36 28.1 39.5 51.1 ReSKGC 81 2023 - - - 39.6 37.3 41.3 43.7 - - - - CD 20 2024 - - - - - - - 37.2 28.8 41 53
Table 11: Experimental statistics under the LP subtask in inductive setting.
Model Year NELL-One Wikidata5M N-Shot MRR Hits@1 Hits@5 Hits@10 N-Shot MRR MR Hits@1 Hits@3 Hits@10 DKRL 36 2016 - - - - - Zero-Shot 23.1 78 5.9 32 54.6 GMatching 41 2018 Five-Shot 20 14 26 31 - - - - - - MetaR 64 2019 Five-Shot 26 17 35 44 - - - - - - GMatching 41 2018 One-Shot 19 12 26 31 - - - - - - RoBERTa 83 2019 - - - - - Zero-Shot 7.4 723 0.7 1 19.6 MetaR 64 2019 One-Shot 25 17 34 40 - - - - - - StAR 88 2021 Zero-Shot 26 17 35 45 - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - KEPLER 47 2021 - - - - - Zero-Shot 40.2 28 22.2 51.4 73 SimKGC 86 2021 - - - - - Zero-Shot 71.4 - 50.9 78.5 91.7 KG-S2S 51 2022 Zero-Shot 31 22 41 48 - - - - - -
Transformer-based models excel in capturing complex patterns and contextual information, significantly enhancing link prediction performance. However, they come with higher computational costs and complexity.
KG-BERT and KGT5, require substantial computational power for training and inference, making them less suitable for environments with limited resources. Models like MTL-KGC and CSPromp-KG involve multi-task learning and prompt engineering, which can complicate the training process and require extensive hyperparameter tuning. While transformer-based models perform well on large-scale datasets, their scalability can be a concern due to the high computational and memory demands. Techniques like advanced negative sampling in SimKGC help mitigate some of these challenges.
5 Future Directions
The integration of large language models (LLMs) into knowledge representation learning (KRL) has demonstrated significant potential in enhancing the capabilities of artificial intelligence systems. However, there are numerous avenues for further research and development. This section outlines six promising directions to advance the field.
5.1 Advanced Integration of LLMs and KRL for Enhanced Contextual Understanding
One major area for future exploration is the advanced integration of LLMs with KRL to achieve deeper contextual understanding. Current models have shown that LLMs can significantly enhance the representation and reasoning capabilities of knowledge graphs (KGs). However, most existing approaches treat textual and structural data separately or integrate them at a superficial level. Future research could focus on more sophisticated methods of merging these data types.
For instance, developing models that can dynamically switch between textual and structural contexts based on the task requirements could lead to more robust and contextually aware systems. This might involve hybrid architectures that combine the strengths of LLMs for textual data with graph neural networks (GNNs) for structural data. These models could leverage attention mechanisms to weigh the importance of textual and structural information dynamically.
Additionally, exploring techniques such as multi-modal learning, where models are trained on diverse types of data (text, images, structured data), could further enhance the understanding and reasoning capabilities of AI systems. By creating a unified representation space for different data types, models could learn to better utilize contextual cues from various sources, leading to more accurate and insightful knowledge representation.
5.2 Development of Efficient Training and Inference Techniques
The computational demands of LLMs are a significant challenge, particularly when integrating them with KRL tasks. Future research should focus on developing more efficient training and inference techniques to make these models more accessible and practical for real-world applications.
Techniques such as model pruning, quantization, and knowledge distillation have shown promise in reducing the computational load of LLMs without significantly sacrificing performance. Applying these techniques specifically to the context of KRL could yield models that are both powerful and efficient. For example, distilling the knowledge from a large LLM into a smaller model that is optimized for KRL tasks could provide a balance between performance and efficiency.
Another promising direction is the use of transfer learning and continual learning approaches. Instead of training models from scratch for each new task or domain, leveraging pre-trained models and fine-tuning them on specific KRL tasks can save significant computational resources. Continual learning techniques, where models incrementally learn from new data without forgetting previously acquired knowledge, can also enhance the adaptability and efficiency of LLM-based KRL systems.
5.3 Improving Generalization and Robustness in KRL Models
While LLM-enhanced KRL models have achieved impressive results on benchmark datasets, their generalization to unseen data and robustness in the face of noisy or incomplete data remain critical challenges. Future research should aim to develop models that can generalize better across diverse domains and handle the inherent uncertainty and noise in real-world data.
One approach to improving generalization is to incorporate more diverse and representative datasets during training. Models trained on a broader range of data are likely to perform better on unseen tasks. Additionally, techniques such as domain adaptation and meta-learning can help models adapt to new domains with minimal additional training.
Robustness can be enhanced by developing models that can handle noisy or incomplete data effectively. This could involve techniques such as data augmentation, adversarial training, and the use of probabilistic graphical models that can explicitly model uncertainty. Ensuring that models can identify and mitigate the impact of noise and handle missing information will be crucial for their deployment in real-world applications.
5.4 Exploring Interpretability and Explainability in KRL Models
As LLM-enhanced KRL models become more complex, their interpretability and explainability become increasingly important, especially in critical applications such as healthcare, finance, and legal domains. Future research should focus on developing methods to make these models more transparent and understandable.
One approach is to develop techniques that can provide insights into the decision-making process of LLM-based KRL models. This could involve the use of attention mechanisms that highlight the most relevant parts of the input data, as well as the development of visualization tools that can illustrate how different parts of the model contribute to the final predictions.
Another promising direction is the integration of explainable AI (XAI) techniques with KRL models. XAI techniques, such as SHAP (SHapley Additive exPlanations) and LIME (Local Interpretable Model-agnostic Explanations), can provide post-hoc explanations for the model’s predictions. Incorporating these techniques into KRL models can help users understand the underlying reasoning behind the model’s outputs, increasing trust and facilitating the adoption of these models in sensitive applications.
5.5 Enhancing Human-AI Collaboration in KRL
The integration of LLMs with KRL offers significant potential for enhancing human-AI collaboration. Future research should explore how these advanced models can be leveraged to support and augment human decision-making and knowledge discovery processes.
Developing interactive systems that allow humans to query and interact with LLM-enhanced KRL models can lead to more effective knowledge exploration and problem-solving. For example, creating natural language interfaces that enable users to ask questions and receive explanations or suggestions from the model can make complex knowledge graphs more accessible and useful.
Additionally, research could focus on collaborative learning frameworks where human feedback is continuously incorporated into the model’s learning process. This could involve techniques such as active learning, where the model identifies uncertain or ambiguous cases and seeks human input to improve its understanding. By fostering a symbiotic relationship between humans and AI, these systems can leverage the strengths of both to achieve superior performance.
5.6 Addressing Ethical and Societal Implications
As LLM-enhanced KRL models become more powerful and pervasive, it is crucial to address the ethical and societal implications of their deployment. Future research should focus on ensuring that these models are developed and used responsibly, with considerations for fairness, accountability, and transparency.
One important aspect is to ensure that these models do not perpetuate or amplify biases present in the training data. Techniques such as bias detection and mitigation, as well as the development of fairness-aware algorithms, can help address these issues. It is also important to consider the broader societal impacts of deploying these models, including potential job displacement and the impact on privacy.
Additionally, research should explore the development of regulatory frameworks and ethical guidelines for the use of LLM-enhanced KRL models. This could involve collaborations between researchers, policymakers, and industry stakeholders to ensure that these models are used in ways that align with societal values and ethical principles.
6 Conclusion
The integration of LLM with KRL represents a transformative advancement in AI, offering significant improvements in the ability to understand, represent, and utilize complex knowledge structures. This survey has systematically reviewed various approaches, categorizing them into encoder-based, encoder-decoder-based, and decoder-based methods, and evaluating their performance across multiple downstream tasks. Finally, there remain critical areas for further research.

### v2

@@
-
+Large Language Model Enhanced Knowledge Representation Learning: A Survey
 Abstract
-The integration of Large Language Models (LLMs) with Knowledge Representation Learning (KRL) signifies a pivotal advancement in the field of artificial intelligence, enhancing the ability to capture and utilize complex knowledge structures.
-This synergy leverages the advanced linguistic and contextual understanding capabilities of LLMs to improve the accuracy, adaptability, and efficacy of KRL, thereby expanding its applications and potential.
-Despite the increasing volume of research focused on embedding LLMs within the domain of knowledge representation, a thorough review that examines the fundamental components and processes of these enhanced models is conspicuously absent.
+The integration of Large Language Models (LLM) with Knowledge Representation Learning (KRL) signifies a significant advancement in the field of artificial intelligence (AI), enhancing the ability to capture and utilize both structure and textual information.
+Despite the increasing research on enhancing KRL with LLMs, a thorough survey that analyse processes of these enhanced models is conspicuously absent.
 Our survey addresses this by categorizing these models based on three distinct Transformer architectures, and by analyzing experimental data from various KRL downstream tasks to evaluate the strengths and weaknesses of each approach.
-Finally, we identify and explore potential future research directions in this emerging yet underexplored domain, proposing pathways for continued progress.
+Finally, we identify and explore potential future research directions in this emerging yet underexplored domain.
 keywords:
-Knowledge graph, Large language model, Representation learning
+Knowledge graph, Large language model, Knowledge representation learning
 1 Introduction
-Large language models (LLMs) (e.g., BERT 18 , LLaMA 59 ) which represents a direction of ever-increasing model sizes pre-trained on larger corpora, have demonstrated powerful capabilities in solving natural language processing (NLP) tasks, including question answering 99 , text generation 100 and document understanding 101 .
-There are no clear and static thresholds regarding the model sizes.
-Early LLMs (e.g., BERT, RoBERTa) adopt an encoder architecture and show capabilities in text representation learning and natural language understanding.
-In recent years, more focus has been given to larger encoder-decoder 102 or decoder-only 103 architectures.
-As the model size scales up, such LLMs have also shown reasoning ability and even more advanced emergent ability 104 , exposing a strong potential for Artificial General Intelligence (AGI).
-This inflection point, with the arrival of LLMs, marks a paradigm shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge.
-As a popular approach for explicit knowledge representation, KGs are now widely investigated for the combination with Transformer-based LLMs, including pre-trained masked language models (PLMs) like BERT and RoBERTa, and more recent generative LLMs like the GPT series and LLaMA.
-Some works use LLMs to augment knowledge graph representation learning.
-In this survey, considering three directions, encoder-based methods, encoder-decoder-based method, and decoder-based methods.
-We present a better understanding of the shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge.
+Large language models (LLM) such as BERT [1], BART [2], and LLaMA [119], represent a trend towards increasingly larger models pre-trained on vast corpora.
+These models have exhibited remarkable capabilities in addressing various natural language processing (NLP) tasks, including text generation [3], sentiment analysis [4], and question answering [5].
+There are no definitive and static thresholds for model sizes.
+Early LLM, like BERT and RoBERTa [6], utilize an encoder architecture and demonstrate proficiency in text representation learning and natural language understanding.
+More recently, there has been a shift towards larger encoder-decoder architectures or decoder architectures.
+As the size of these models grows, they exhibit advanced reasoning abilities and complex task-solving capabilities [7], indicating significant potential for a wide range of AI applications.
+This inflection point, catalyzed by the introduction of LLM, heralds a paradigm shift from structured knowledge representation to the integration of structured and textual representations.
+Knowledge Graphs (KG), once primarily used for structured knowledge representation, are now being extensively explored in conjunction with Transformer-based LLM, such as pre-trained language models (PLMs) and recent generative LLM.
+Various studies leverage LLM to refine the processes of knowledge representation learning (KRL).
+In this survey, we categorize the approaches into three distinct methodologies: encoder-based, encoder-decoder-based, and decoder-based methods.
+This categorization aids in understanding the shift from purely structured representation to a hybrid model that effectively combines structured and textual knowledge representations.
+There are few existing surveys that investigate the enhancement of KRL by LLM.
 Cao et al.
-22 and Biswas et al.
-40 discuss recent advancements in knowledge graph representation learning, yet they inadequately address aspects related to the integration with large models.
+[8] and Biswas et al.
+[9] discuss recent advancements in KRL, yet they inadequately address aspects related to the integration with LLM.
 Pan et al.
-42 and Pan et al.
-43 explore the combination of knowledge graphs and large models, specifically addressing LLM4KG and KG4LLM; however, they provide limited coverage of representation learning.
-Consequently, there is currently no comprehensive review article dedicated to outlining the latest developments in the field of knowledge graph representation learning.
-Contributions.   The notable contributions of this survey are summarized as follows:
+[10] and Pan et al.
+[11] explore the combination of KG and LLM, specifically addressing LLM4KG and KG4LLM.
+However, they provide limited coverage of KRL.
+Consequently, there is currently no comprehensive review article dedicated to outlining the latest developments in the field of KRL.
+Contributions. The contributions of this survey can be summarized as follows:
+  • LLM Enhanced KRL Survey.
+This work provides the first comprehensive survey on the LLM enhanced KRL.
+Specifically, we analyze how different encoder-based, encoder-decoder-based, and decoder-based methods have been enhanced through the integration with LLM, offering a new dimension to understanding their capabilities and limitations in KRL.
   • Categorization of KRL.
-We systematically summarize the knowledge representation learning where large language models can be adopted into: encoder-based, encoder-decoder-based, and decoder-based methods.
-  • Systematic Review of Techniques.
-We provide the most comprehensive overview of large language models on knowledge graph representation learning techniques.
-For different methods, we summarize the representative models, provide detailed illustrations, and make necessary comparisons.
-  • Future Directions. We delve into the foundational principles of language models on graphs and propose six prospective avenues for future exploration.
+We categorize the existing methods in KRL according to their foundational LLM architectures, providing detailed descriptions and comparisons of these methods from the perspective of their computational models and the enhancement strategies employed.
+  • Future Directions.
+We propose some directions for future research in KRL by highlighting unique properties of LLM enhanced methods.
+These insights are expected to inspire advancements not only within the realm of KRL but also in broader fields like NLP, leveraging the robust capabilities of LLM.
 Organization.
 The structure of this survey is organized as follows.
-Section 2 introduce the foundation by introducing the foundational concepts pertinent to this review, including Knowledge Graphs, Large Language Models, and Text-based Representation Learning.
-Sections 3 investigates the core methodologies employed in current research, categorizing them into three types of Transformer, respectively.
+Section 2 (§2) introduces the foundational concepts that are essential to this survey, including Knowledge Graph, Large Language Model, and Knowledge Representation Learning.
+Sections 3 (§3) investigates the core methodologies employed in current research, categorizing them into three types of Transformer, respectively.
 Each of these sections further explores various sub-methodologies.
-In Section 4, we discuss the advantages of each type of approach from the performance evaluation results of each downstream task.
-Section 5 outlines potential future research directions that align with the discussed categories, proposing areas that promise significant advancements.
-The conclusion in Section 6 synthesizes the insights drawn from the review and highlights the implications for future research.
+In Section 4 (§4), we discuss the advantages of each type of approach from the performance evaluation results of each downstream task.
+Section 5 (§5) outlines potential future directions that align with the discussed categories.
+The conclusion in Section 6 (§6) provides a comprehensive summary of the survey.
 2 Preliminaries
@@
 One of the significant advantages of pretraining in NLP is that training data can be derived from any unlabeled text corpus, providing an essentially unlimited amount of training data.
-Early pretraining methods in NLP, such as the Neural Language Model (NLM) and Word2vec, were static, meaning they could not adapt effectively to different semantic environments.
+Early pretraining methods in NLP, such as the Neural Language Model (NLM) [12] and Word2vec [13], were static, meaning they could not adapt effectively to different semantic environments.
 Static models learn fixed word representations, which limits their ability to handle context variations.
-To overcome this limitation, dynamic pretraining techniques were developed, leading to Pretrained Language Model (PLM) like BERT (Bidirectional Encoder Representations from Transformers) and XLNet.
+To overcome this limitation, dynamic pretraining techniques were developed, leading to Pretrained Language Model (PLM) like BERT (Bidirectional Encoder Representations from Transformers) and XLNet [14].
 These models dynamically adjust word representations based on context, significantly enhancing their adaptability and performance across various tasks.
@@
 where $Q,K,V\in\mathds{R}^{N_{S}\times d_{k}}$ are the query, key, and value vectors for each word in the sentence, respectively.
-Following BERT, other masked language models are proposed, such as RoBERTa, ALBERT, and ELECTRA, with similar architectures and objectives of text representation.
-Although the original Transformer paper was experimented on machine translation, it was not until the release of GPT-2 that language generation (aka.
+Following BERT, other masked language models are proposed, such as RoBERTa, ALBERT [15], and ELECTRA [16], with similar architectures and objectives of text representation.
+Although the original Transformer paper was experimented on machine translation, it was not until the release of GPT-2 [17] that language generation (aka.
 causal language modeling) became impactful on downstream tasks.
@@
 GPT-3.5 was trained on a diverse blend of text and code, making it adept at a wide range of tasks from conversation to code generation.
-ChatGPT leverages reinforcement learning from human feedback (RLHF), which has emerged as a promising approach to align large language models with human intent.
+ChatGPT leverages reinforcement learning from human feedback (RLHF) [18], which has emerged as a promising approach to align large language models with human intent.
 This alignment is achieved through iterative tuning based on human feedback, ensuring the responses are more relevant and accurate.
 Figure 1: The history and evolution of LLM.
-We would like to point out that the word “large” in LLM is not associated with a clear and static threshold to divide language models.
-“Large” actually refers to a direction in which language models are inevitably evolving, and larger foundational models tend to possess significantly more representation and generalization power.
-Hence, we define LLMs to encompass both medium-scale PLMs, such as BERT, and large-scale LMs, like GPT-4, as suggested.
+We would like to point out that the word ‘large’ in LLM is not associated with a clear and static threshold to divide language models.
+‘Large’ actually refers to a direction in which language models are inevitably evolving, and larger foundational models tend to possess significantly more representation and generalization power.
+Hence, we define LLM to encompass both medium-scale PLMs, such as BERT, and large-scale LMs, like GPT-4, as suggested.
 2.3 Knowledge Representation Learning
@@
 Convolutional neural networks (CNNs) are effective for encoding entity descriptions and textual mentions, learning deep expressive features from textual information.
-DKRL(CNN) assumes that word order in entity descriptions involves implicit relations between entities that KGs may omit.
+DKRL(CNN) assumes that word order in entity descriptions involves implicit relations between entities that KG may omit.
 It uses a five-layer CNN to discover these implicit relations.
@@
 Table 1: Summary of the embeddings, extended models and combination mechanism in text-based KG embedding.
-Models Structual embedding Textual embedding Extended model Combination mechanism WVV $r_{s}\in\mathbb{R}^{d}$ $h^{(r_{s})}_{d},t^{(r_{s})}_{d}\in\mathbb{R}^{d}$ SE /TransE/ $f_{r}(h,t)=-\|h^{(r_{s})}_{d}-r_{s}+t^{(r_{s})}_{d}\|$ DKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE RTKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $r_{d}\in\mathbb{R}^{d}$ TransE $f_{r}(h,t)=f_{s}+f_{d}$ RDRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE STKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE Joint $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE $h=g_{h}\otimes h_{s}+(1-g_{h})\otimes h_{d}$ $t=g_{t}\otimes t_{s}+(1-g_{t})\otimes t_{d}$ $r=r_{s}$ EDGE $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
+Models Structual embedding Textual embedding Extended model Combination mechanism
+WVV $r_{s}\in\mathbb{R}^{d}$ $h^{(r_{s})}_{d},t^{(r_{s})}_{d}\in\mathbb{R}^{d}$ SE /TransE/ $f_{r}(h,t)=-\|h^{(r_{s})}_{d}-r_{s}+t^{(r_{s})}_{d}\|$
+DKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
+RTKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $r_{d}\in\mathbb{R}^{d}$ TransE $f_{r}(h,t)=f_{s}+f_{d}$
+RDRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
+STKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
+Joint $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE $h=g_{h}\otimes h_{s}+(1-g_{h})\otimes h_{d}$
+$t=g_{t}\otimes t_{s}+(1-g_{t})\otimes t_{d}$
+$r=r_{s}$
+EDGE $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
 Table 2: Notations of Concepts.
-[…]
+Notation Description
+$h$, $r$, $t$ Head entity, Relation, Tail entity
+$\mathbf{X}{(h,r)}$, $\mathbf{X}(t)$ Input sequences for head-relation and tail entity
+$D$, $D^{+}$, $D^{-}$ Set of all triples, positive triples, negative triples
+$\mathbf{u}$, $\mathbf{v}$ Contextualized embeddings of head-relation pair and tail entity
+$s_{c}$, $s_{d}$ Classification score, Distance score
+$\mathbf{h}$, $\mathbf{r}$, $\mathbf{t}$ Embeddings for head entity, relation, and tail entity
+$L_{c}$, $L_{d}$, $L$ Classification loss, Distance-based loss, Total loss
+$\lambda$, $\gamma$, $n$, $K$ Margin, Weight factor, Number of negative samples, Sum limit
+$\sigma$ Sigmoid function
+$\alpha$, $\beta$ Auxiliary loss weights
+$\mathbf{K}$, $I$, $X$ Transformed structural embeddings, Instruction prompt, Triple prompt
 Table 3: An overview of various LLM enhanced KRL models.
-Model Type Base Model Source Code 2019 ERNIE Encoder BERT https://github.com/thunlp/ERNIE KnowBert Encoder BERT https://github.com/allenai/kb KG-BERT Encoder BERT https://github.com/yao8839836/kg-bert 2020 MTL-KGC Encoder BERT https://github.com/bosung/MTL-KGC Pretrain-KGE Encoder BERT - K-BERT Encoder BERT https://github.com/autoliuweijie/K-BERT 2021 StAR Encoder BERT, RoBERTa https://github.com/wangbo9719/StAR_KGC KG-GPT2 Decoder GPT2 - MEM-KGC Encoder BERT - LaSS Encoder BERT, RoBERTa https://github.com/jhshen95/LASS KEPLER Encoder RoBERTa https://github.com/THU-KEG/KEPLER BLP Encoder BERT https://github.com/dfdazac/blp SimKGC Encoder BERT https://github.com/intfloat/SimKGC 2022 LP-BERT Encoder BERT, RoBERTa - PKGC Encoder BERT, RoBERTa, LUKE https://github.com/THU-KEG/PKGC Nayyeri et al.
-Encoder BERT https://github.com/ZihaoWang/Hypercomplex-KG-Embedding CoDEx Encoder Sentence-BERT - KGT5 Encoder-Decoder T5 https://github.com/apoorvumang/kgt5 OpenWorld KGC Encoder BERT - Lovelace et al.
-Encoder BERT, PubMedBERT https://github.com/justinlovelace/LM-KG-Completion kNN-KRL Encoder BERT https://github.com/zjunlp/KNN-KG LMKE Encoder BERT https://github.com/Neph0s/LMKE GenKGC Encoder-Decoder BART https://github.com/zjunlp/PromptKG/tree/main/research/GenKGC KG-S2S Encoder-Decoder T5 https://github.com/chenchens190009/KG-S2S 2023 LambdaKG Encoder-Decoder BERT, BART, T5 https://github.com/zjunlp/PromptKG/tree/main/lambdaKG CSPromp-KG Encoder BERT https://github.com/chenchens190009/CSProm-KG ReSKGC Encoder-Decoder T5 - StructGPT Decoder GPT 3.5 https://github.com/RUCAIBox/StructGPT ReasoningLM Encoder RoBERTa https://github.com/RUCAIBox/ReasoningLM KG-GPT Decoder GPT 3.5 https://github.com/jiho283/KG-GPT Keqing Decoder GPT 3.5, LLaMA 2 - AutoKG Decoder GPT 3.5 https://github.com/zjunlp/AutoKG KG-LLM Decoder ChatGLM, LLaMA 2 https://github.com/yao8839836/kg-llm KoPA Decoder Alpaca, LLaMA, GPT 3.5 https://github.com/zjukg/KoPA 2024 CD Decoder PaLM2 https://github.com/DavidLi0406/Contextulization-Distillation CP-KGC Decoder Qwen, LLaMA 2, GPT 4 https://github.com/sjlmg/CP-KGC KICGPT Decoder GPT 3.5 - KnowledgeNavigator Decoder GPT 3.5, LLaMA 2 -
-Figure 2: A taxonomy of LLM enhanced KRL.
+Year Model Type Base Model Source Code
+2019 KG-BERT Encoder BERT Github
+2020 MTL-KGC Encoder BERT Github
+Pretrain-KGE Encoder BERT -
+K-BERT Encoder BERT Github
+2021 StAR Encoder BERT, RoBERTa Github
+KG-GPT2 Decoder GPT2 -
+BERT-ResNet Encoder BERT Github
+MEM-KGC Encoder BERT -
+LaSS Encoder BERT, RoBERTa Github
+KEPLER Encoder RoBERTa Github
+BLP Encoder BERT Github
+SimKGC Encoder BERT Github
+MLMLM Encoder RoBERTa Github
+2022 LP-BERT Encoder BERT, RoBERTa -
+PKGC Encoder BERT, RoBERTa, LUKE Github
+KGT5 Encoder-Decoder T5 Github
+OpenWorld KGC Encoder BERT -
+Lovelace et al. Encoder BERT, PubMedBERT Github
+$k$NN-KGE Encoder BERT Github
+LMKE Encoder BERT Github
+GenKGC Encoder-Decoder BART Github
+KG-S2S Encoder-Decoder T5 Github
+2023 LambdaKG Encoder-Decoder BERT, BART, T5 Github
+CSPromp-KG Encoder BERT Github
+ReSKGC Encoder-Decoder T5 -
+ReasoningLM Encoder RoBERTa Github
+KG-LLM Decoder ChatGLM, LLaMA 2 Github
+KoPA Decoder Alpaca, LLaMA, GPT 3.5 Github
+2024 CD Decoder PaLM2 Github
+CP-KGC Decoder Qwen, LLaMA 2, GPT 4 Github
+KICGPT Decoder GPT 3.5 -
 3 LLM Enhanced KRL Methods
-In this section, we delve into the advanced methodologies for enhancing Knowledge Representation Learning using large language models.
-The section is structured into three main sections, each focusing on a different category of LLM-based methods.
-Section 3.1 explores Encoder-based methods, which utilize powerful encoders to generate dense vector representations of knowledge entities.
-Section 3.2 discusses Encoder-Decoder-based methods that leverage both encoding and decoding processes to capture and generate complex relations within knowledge graphs.
-Finally, Section 3.3 examines Decoder-based methods, emphasizing the generation and prediction capabilities of decoders to infer and expand knowledge representations.
+LLM have significantly advanced the field of KRL.
+These models leverage vast amounts of data and sophisticated architectures to generate context-aware representations of knowledge.
+This section explores various methods for enhancing KRL using LLM, categorized into three main approaches: encoder-based methods (§3.1), encoder-decoder-based methods (§3.2), and decoder-based methods (§3.3).
+Each category leverages different aspects of LLM to optimize the encoding, decoding, and overall integration of knowledge graph data, ultimately improving the accuracy and utility of knowledge representations in diverse applications.
+The evolution of these models is illustrated in Fig.
+2.
 Figure 2: A flowchart of model evolution.
 3.1 Encoder-Based Methods
-These works employ the encoder of Transformer to transform knowledge graph elements into textual sequences that are then processed to evaluate triple plausibility or to complete knowledge graphs.
-The contextual capabilities are central to assessing relations and entity attributes in the graph structure.
-3.1.1 Multi-Level Encoder
-The typical approach is illustrated in Fig.
-4.
-This category primarily focuses on employing BERT or similar encoder-based pre-trained language models to convert knowledge graph elements into textual sequences, which are then used to assess the plausibility of triples.
-The common method is to use BERT deep contextual embeddings to enhance the representation of entities and relations, thereby improving accuracy in predicting plausible triples.
-Figure 4: An overview of multi-level encoder method.
-$\text{Input}=[CLS]+\text{Tokenize}(h)+[SEP]+\text{Tokenize}(r)+[SEP]+\text{Tokenize}(t)+[SEP]$ (4)
-where [CLS] and [SEP] are special tokens used by BERT, and $\text{Tokenize}(\cdot)$ represents the tokenization process.
-The classification task is to determine if the concatenated sequence correctly represents a valid triple from the KG. A logistic sigmoid function computes the probability of the triple being valid:
-$s_{\tau}=\sigma(\mathbf{C}\mathbf{W}^{T})$ (5)
-Here, $\mathbf{C}$ is the final hidden state of the [CLS] token, $\mathbf{W}$ is a learnable parameter matrix, and $\sigma$ denotes the sigmoid function.
-The training uses a cross-entropy loss function to handle both positive and negative examples of triples:
-$L=-\sum_{\tau\in D^{+}\cup D^{-}}\left(y_{\tau}\log(s_{\tau,0})+(1-y_{\tau})\log(s_{\tau,1})\right)$ (6)
-where $D^{+}$ and $D^{-}$ represent the sets of positive and negative triples, respectively, and $y_{\tau}$ is the label (0 or 1) of the triple.
-The use of BERT allows the model to leverage contextualized text embeddings, improving the quality of predictions compared to traditional methods that might only use structural information of the KG.
-Pretrain-KGE 74 enriches traditional Knowledge Graph Embeddings (KRLs) using BERT pre-trained embeddings, aiming to leverage linguistic context for better entity and relation representation.
-The method is beneficial in contexts where semantic richness is critical, but it relies heavily on the quality and scope of the training data.
-StAR 88 introduces a hybrid model combining textual encoding with graph embedding techniques, addressing the limitations of both approaches by modeling graph elements and their spatial relations.
-It benefits from enhanced representational efficiency but might struggle with complexity and scalability due to the dual nature of the encoding.
-MEM-KGC 63 utilizes a Masked Language Model approach for predicting masked entities in a knowledge graph, enhancing embeddings for better performance on KGC tasks.
-Its strength lies in handling unseen entities during testing, though it might introduce biases from the pre-training data used in the MLM.
-The advantage of this approach lies in leveraging the robust linguistic capabilities of BERT, which can capture complex patterns in text data that are beneficial for knowledge graph completion.
-However, the reliance on textual representation may limit the model’s ability to capture purely structural relationships inherent in knowledge graphs, potentially leading to performance bottlenecks when dealing with entities that have poor textual descriptions.
-3.1.2 Score Function Variation
-This category focuses on utilizing pre-trained language models across multiple tasks simultaneously, such as link prediction, entity resolution, and relation classification.
-Techniques often involve fine-tuning a shared language model on KG-specific data, employing strategies that allow the model to leverage learned representations more effectively.
-The process of the typical method is depicted in Fig.
+Encoder-based methods primarily utilize encoders such as BERT [1] and RoBERTa [6].
+These methods are distinguished based on how they encode KG triples.
+In triple encoding (§3.1.1), the head entity, relation, and tail entity are encoded together as a single unit.
+Translation encoding (§3.1.2), on the other hand, encodes the head entity and relation together, while the tail entity is encoded separately, with optimization performed using vector space distances.
+Independent encoding (§3.1.3) treats the head entity, relation, and tail entity as separate inputs, encoding each independently.
+3.1.1 Triple Encoding
+Triple encoding methods aim to represent the entire KG triple as a single unit.
+This approach leverages the combined semantic and relational information of the head entity, relation, and tail entity, providing a holistic representation that can enhance downstream tasks.
+An example of a typical model KG-BERT [19], that using triple encoding is illustrated in Fig.
+3.
+Figure 3: An overview of triple encoding method.
+It treats triples as textual sequences and fine-tunes a pre-trained BERT model to predict the plausibility of a given triple.
+The model takes as input the descriptions of the head entity $h$, relation $r$, and tail entity $t$:
+[…]
+where $\text{Tok}_{hi}$, $\text{Tok}_{ri}$, and $\text{Tok}_{ti}$ represent the tokens of the head entity, relation, and tail entity descriptions, respectively.
+The model computes the final hidden vector of the special [CLS] token, which serves as the aggregate sequence representation.
+The plausibility score of a triple $\tau=(h,r,t)$ is then computed as follows:
+$s_{\tau}=\text{sigmoid}(CW^{T})$ (5)
+where $C$ is the final hidden state corresponding to the [CLS] token, and $W$ is the classification layer weight.
+The loss function for training is the cross-entropy loss computed over the positive and negative triples:
+$L=-\sum_{\tau\in D^{+}\cup D^{-}}(y_{\tau}\log(s_{\tau 0})+(1-y_{\tau})\log(s_{\tau 1}))$ (6)
+where $y_{\tau}\in\{0,1\}$ is the label indicating whether the triple is positive or negative, and $D^{+}$ and $D^{-}$ are the sets of positive and negative triples, respectively.
+The negative triples are generated by corrupting either the head or the tail entity of a positive triple.
+MTL-KGC [20] utilizes a multi-task learning framework to enhance knowledge graph completion.
+It incorporates additional tasks like relation prediction and relevance ranking, significantly improving upon the limitations of prior models such as KG-BERT by learning more relational properties and distinguishing between lexically similar candidates effectively.
+K-BERT [21] integrates structured domain knowledge from knowledge graphs directly into the pre-training process.
+It maintains the original BERT architecture but adds a mechanism to inject relevant triples from a knowledge graph into the training examples.
+This approach aims to enrich the language model with domain-specific information, addressing BERT’s limitation in handling domain-specific tasks without extensive retraining.
+MLMLM [22] introduces a novel method for link prediction by employing Mean Likelihood Masked Language Models to directly generate potential entities.
+This approach allows for leveraging the knowledge embedded in pre-trained MLMs to enhance the interpretability and scalability of knowledge graphs.
+PKGC [23] is a new PLM-based model for KGC that utilizes prompt engineering to enhance knowledge utilization from PLMs.
+This model adapts PLMs for the KGC task by converting knowledge triples into natural language prompts, thereby facilitating the effective use of PLMs’ latent knowledge.
+CSProm-KG [24] innovatively incorporates both structural and textual information into the knowledge graph completion process.
+This model employs Conditional Soft Prompts that adapt the inputs based on the structure of the knowledge graph, allowing pre-trained language models to efficiently utilize both types of information.
+3.1.2 Translation Encoding
+Translation encoding methods focus on encoding the head entity and relation together while treating the tail entity separately.
+This technique relies on optimizing the vector space distances between these entities, facilitating more accurate knowledge representation and reasoning.
+Fig.
+4 depicts a representative model StAR [25] that employs translation encoding.
+Figure 4: An overview of translation encoding method.
+This model partitions each triple into two asymmetric parts: one combining the head entity and relation, and the other with the tail entity.
+The StAR model then uses a Siamese-style textual encoder to encode these parts into contextualized representations.
+The translation function is formulated as follows:
+$\mathbf{u}=\text{Pool}(\text{Transformer-Enc}(\mathbf{X}_{(h,r)})),\quad\mathbf{v}=\text{Pool}(\text{Transformer-Enc}(\mathbf{X}_{t}))$ (7)
+where $\mathbf{X}_{(h,r)}=[\text{CLS}]\;\mathbf{X}_{h}\;[\text{SEP}]\;\mathbf{X}_{r}\;[\text{SEP}]$ and $\mathbf{X}_{t}=[\text{CLS}]\;\mathbf{X}_{t}\;[\text{SEP}]$ are the input sequences for the head-relation pair and the tail entity, respectively.
+$\text{Pool}(\cdot)$ denotes the pooling operation to obtain the sequence-level representation from the transformer’s output.
+The plausibility score of a triple $(h,r,t)$ is then determined by a combination of a deterministic classifier and a spatial measurement:
+$s_{c}=\text{softmax}(\text{MLP}([\mathbf{u};\mathbf{u}\circ\mathbf{v};\mathbf{u}-\mathbf{v};\mathbf{v}])),\quad s_{d}=-\|\mathbf{u}-\mathbf{v}\|_{2}$ (8)
+where $[\mathbf{u};\mathbf{u}\circ\mathbf{v};\mathbf{u}-\mathbf{v};\mathbf{v}]$ represents the interactive concatenation of the contextualized embeddings $\mathbf{u}$ and $\mathbf{v}$, and $\text{MLP}(\cdot)$ is a multi-layer perceptron.
+The model is trained using both a triple classification objective and a contrastive loss:
+$L_{c}=-\frac{1}{|D|}\sum_{t_{p}\in D}\left(\log s_{c}+\sum_{t_{p}^{\prime}\in N(t_{p})}\log(1-s_{c}^{\prime})\right)$ (9)
+$L_{d}=\frac{1}{|D|}\sum_{t_{p}\in D}\sum_{t_{p}^{\prime}\in N(t_{p})}\max(0,\lambda-s_{d}+s_{d}^{\prime})$ (10)
+where $D$ is the set of positive triples, $N(t_{p})$ denotes the set of negative triples generated by corrupting positive triples, and $\lambda$ is the margin.
+The overall training loss is a weighted sum of the two objectives:
+$L=L_{c}+\gamma L_{d}$ (11)
+where $\gamma$ is a hyperparameter controlling the balance between the two losses.
+SimKGC [26] utilizes a bi-encoder architecture to improve contrastive learning efficiency in knowledge graph completion.
+It integrates three types of negative sampling, in-batch, pre-batch, and self-negatives, enhancing training effectiveness by leveraging larger negative sets and a novel InfoNCE loss function.
+LP-BERT [27] employs a multi-task pre-training strategy, enhancing knowledge graph completion by predicting linkage between entities.
+This method integrates masked language, entity, and relation modeling tasks, significantly improving link prediction.
+It effectively handles unseen entities and relations by leveraging contextualized entity and relation information from large datasets.
+3.1.3 Independent Encoding
+Independent encoding methods separately encode each component of a triple: the head entity, relation, and tail entity.
+This approach allows for flexible and modular representations, which can be beneficial for certain types of knowledge graph applications.
+The architecture of a typical independent encoding model KEPLER [28] is shown in Fig.
 5.
-Figure 5: An overview of score function variation method.
-Papers in this subcategory explore the enhancement of knowledge graph embeddings through multi-task learning frameworks and fine-tuning processes.
-These approaches often integrate additional tasks like relation prediction and relevance ranking with the primary link prediction task, aiming to create a more holistic learning scenario that captures various aspects of knowledge graphs.
-Each KG triplet $(h,r,t)$ is represented by natural language descriptions of the head $h$, relation $r$, and tail $t$.
-These descriptions are tokenized into sequences $T_{h}$, $T_{r}$, and $T_{t}$.
-A pre-trained language model (LM) encodes these sequences.
-The output embeddings for the head, relation, and tail are obtained by mean pooling the LM outputs for respective tokens.
-The semantic embedding process uses the formula:
-$h=\text{MeanPool}(\text{LM}(T_{h}))$ (7)
-$r=\text{MeanPool}(\text{LM}(T_{r}))$ (8)
-$t=\text{MeanPool}(\text{LM}(T_{t}))$ (9)
-where $T_{h},T_{r},T_{t}$ are token sequences for the head, relation, and tail, respectively.
-After semantic embedding, a structured loss is optimized to align the embeddings with the KG structure.
-This loss is designed to make the embedding of a true triplet $(h,r,t)$ such that $h+r$ approximates $t$.
-The structure loss is defined by the scoring function:
-$f(h,r,t)=b-\frac{1}{2}\|h+r-t\|^{2}$ (10)
-where $b$ is a margin parameter. The loss aims to minimize the Euclidean distance between the predicted and actual tail entity in the embedding space.
-The model estimates the probability of a triplet being correct based on the structured score function.
-To efficiently handle large KGs, negative sampling is employed, which involves generating negative triplets by corrupting either the head or the tail entity.
-MTL-KGC 66 integrates multiple learning tasks (link prediction, relation prediction, relevance ranking) to enhance KGC using a shared BERT layer.
-Multi-task learning helps in leveraging linguistic information across different tasks, enhancing overall learning efficiency.
-However, it may suffer from task interference where one tasks learning negatively impacts another.
-SimKGC 86 introduces an efficient contrastive learning strategy with advanced negative sampling for text-based KGC.
-It focuses on hard negatives to enhance model training, which improves learning efficiency but may need extensive tuning to optimize the negative sampling strategy.
-KEPLER 47 fuses the capabilities of PLMs with traditional KE techniques, offering dual optimization on both factual and linguistic fronts.
-This integration allows for richer representations, though the complexity of joint optimization could lead to challenges in model convergence.
-The multi-task learning approach helps in capturing a more comprehensive semantic understanding by tackling multiple aspects of knowledge graph completion simultaneously.
-This can lead to models that generalize better on unseen data.
-However, these methods can be computationally expensive and complex to tune, as they require balancing multiple learning objectives which might conflict.
-3.1.3 Negative Strategy
-As shown in Fig.
-6, some integrate textual information encoded by LLMs with structured knowledge graph embeddings.
-This fusion approach enhances the ability to handle both the semantic nuances captured through text and the explicit relational data represented in knowledge graphs.
-Figure 6: An overview of negative strategy method.
-Research in this area focuses on hybrid models that combine the strengths of both textual and structural data.
-Methods typically involve using language models for initial encoding followed by integration with graph embedding techniques, which help in capturing both textual nuances and relational structures within the graph.
-Each entity $v$ and relation $e$ in the graph $\mathcal{G}$ is embedded into a $4\mathbf{D}$-dimensional hypercomplex space, capturing both the structural and textual information:
-$v=\mathbf{s}_{v}+\mathbf{x}_{v_{1}}\mathbf{i}+\mathbf{y}_{v_{2}}\mathbf{j}+\mathbf{z}_{v_{3}}\mathbf{k}$ (11)
-$e=\mathbf{s}_{e}+\mathbf{x}_{e_{1}}\mathbf{i}+\mathbf{y}_{e_{2}}\mathbf{j}+\mathbf{z}_{e_{3}}\mathbf{k}$ (12)
-where $\mathbf{s}_{v}$ and $\mathbf{s}_{e}$ represent the structural embeddings, and $\mathbf{x}_{v_{1}},\mathbf{y}_{v_{2}},\mathbf{z}_{v_{3}}$, etc., are embeddings derived from textual information.
-The compatibility of a triplet $(v,e,w)$ in $\mathcal{G}$ is scored using a function that measures the alignment of the Dihedron product of the embeddings:
-$f(v,e,w)=\mathbf{Re}\left((\overline{v\otimes e})\otimes w\right)$ (13)
-where $\otimes$ denotes the Dihedron product, and $\overline{v\otimes e}$ represents the conjugate of the product.
-The method stands out by efficiently integrating diverse information sources and leveraging the mathematical properties of hypercomplex spaces to enhance the semantic richness of knowledge graph embeddings.
-CoDEx 28 guides traditional KRL models using embeddings derived from LLMs, enhancing the link prediction capabilities.
-It benefits from the textual richness of LLMs but depends heavily on the initial quality of the language model embeddings.
-OpenWorld KGC 72 presents a unified learning framework that combines MLM with multi-task learning to integrate new entities into KGs effectively.
-It excels in handling unconnected entities but might struggle with data sparsity issues.
-Lovelace et al.
-61 optimizes PLMs for KGC by decoupling querying and candidate retrieval embeddings, using supervised and unsupervised techniques.
-It enhances embedding suitability for KGC, but maintaining scalability and efficiency can be challenging.
-The integration of textual and structural embeddings allows for a richer representation of knowledge graphs, potentially improving accuracy in tasks such as link prediction and entity resolution.
-Nevertheless, the complexity of these models can lead to challenges in training stability and increased computational demands.
+Figure 5: An overview of independent encoding method.
+The KEPLER model employs the following methodology for entity descriptions as embeddings:
+$h=E_{\langle s\rangle}(\text{texth}),\quad t=E_{\langle s\rangle}(\text{textt}),\quad r=T_{r},$ (12)
+where texth and textt are the descriptions for the head and tail entities, respectively, with a special token $\langle s\rangle$ at the beginning.
+$T\in\mathbb{R}^{|R|\times d}$ is the relation embeddings matrix, and $h$, $t$, and $r$ are the embeddings for the head entity, tail entity, and relation.
+The knowledge embedding objective of KEPLER uses negative sampling for optimization:
+$L_{KE}=-\log\sigma(\gamma-d_{r}(h,t))-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(d_{r}(h^{\prime}_{i},t^{\prime}_{i})-\gamma),$ (13)
+where $(h^{\prime}_{i},r,t^{\prime}_{i})$ are negative samples, $\gamma$ is the margin, $\sigma$ is the sigmoid function, and $d_{r}$ is the scoring function:
+$d_{r}(h,t)=\|h+r-t\|_{p},$ (14)
+with the norm $p$ set to 1. The negative sampling policy fixes the head entity and randomly samples a tail entity, and vice versa.
+The model BERT-ResNet [29] extends the capabilities of encoder-based KRL methods by combining BERT with a deep residual network, improving the handling of sparse connectivity in knowledge graphs.
+This model leverages the robustness of BERT embeddings and the expressiveness of ResNet’s deep convolutional architecture, significantly enhancing entity ranking performance even with limited training data.
+BLP [30] model focuses on inductive link prediction by utilizing textual entity descriptions and pre-trained language models, emphasizing its ability to generalize well across unseen entities.
+This model outperforms previous approaches by incorporating dynamic graph embeddings, allowing it to effectively adapt to continuously evolving knowledge graphs without retraining.
 3.2 Encoder-Decoder-Based Methods
-In this category, LLMs are employed as sequence-to-sequence generators in knowledge graph completion.
-The models receive textual sequences of query triples and directly generate the text of the tail entity, employing the full generative capabilities of language models to predict missing elements of the graph.
-3.2.1 Multi-Level Encoder
-Recent works explore advanced contextual embedding strategies and contrastive learning methods to refine knowledge graph completion techniques.
-The use of rich contextual embeddings from language models combined with innovative training techniques like contrastive loss helps in distinguishing between correct and incorrect triples more effectively.
+Encoder-decoder-based methods utilize models such as BART [2] and T5 [31].
+These methods are categorized based on the type of input sequence used by the model.
+Structural Encoding methods (§3.2.1) are similar to Encoder-based methods and use encoder-type triple sequences as input.
+Fine-tuning methods (§3.2.2), similar to Decoder-based methods, use natural language expressions of triples as input for the decoder.
+3.2.1 Structural Encoding
+Structural encoding leverages the structure of the triples, feeding them into the encoder as sequences.
+This method captures the syntactic and relational structure of the triples, enhancing the overall representation learned by the model.
 Fig.
-7 outlines the steps involved in the conventional method.
-Figure 7: An overview of multi-level encoder method.
-This subgroup applies contextual embeddings and contrastive learning techniques to knowledge graph completion.
-Contextual methods leverage language models to generate rich, nuanced embeddings, while contrastive approaches focus on distinguishing between correct and incorrect triples through advanced sampling strategies and loss functions.
-For discrimination-based methods, the input sequence for a KG triplet $(v,e,w)$ is prepared as:
-$\mathbf{X}_{\text{pair}}=[CLS]\,\mathbf{x}_{v}[SEP]\mathbf{x}_{e}[SEP]$ (14)
-$\mathbf{X}_{\text{tail}}=[CLS]\,\mathbf{x}_{w}[SEP]$ (15)
-where $\mathbf{x}_{v},\mathbf{x}_{e},\mathbf{x}_{w}$ are the text representations of the head entity, relation, and tail entity, respectively.
-For generation-based methods, the scoring function for a KG triplet is defined using the output probabilities of the sequence model:
-This represents the likelihood of generating the tail entity sequence given the head entity and relation sequences.
-This method represents a significant step forward in KRL research by combining the strengths of PLMs with graph representation learning, thereby improving the effectiveness of knowledge extraction and reasoning tasks.
-KG-GPT2 49 employs GPT-2 for KGC, treating triples as sentences for sequence classification.
-It leverages GPT-2’s contextual understanding but may not always capture the structural nuances of KGs effectively.
-GenKGC 40 transitions KGC to a Seq2Seq generation problem using PLMs, focusing on efficient entity generation. It reduces inference times but may sacrifice accuracy for speed.
-CD 20 distills context-rich information from LLMs into KGC models, enhancing them with detailed text-based contexts.
-It improves contextual understanding but requires significant computational resources for the distillation process.
-Contextual methods benefit from the deep semantic understanding provided by pre-trained language models, which can significantly enhance the interpretability and performance on complex triples.
-Contrastive methods improve learning efficiency by focusing on hard negatives.
-However, these techniques can be sensitive to the choice of negatives and require careful tuning of the contrastive loss parameters.
-3.2.2 Soft Prompt
-This group utilizes sequence-to-sequence and Transformer architectures to perform tasks such as predicting missing entities in triples or completing parts of the knowledge graph.
-The flexibility of Seq2Seq models is particularly useful for tasks where structured inputs need to be transformed into textual or other structured outputs.
-The sequence of actions in the typical method is demonstrated in Fig.
-8.
-Figure 8: An overview of soft prompt method.
-Seq2Seq and Transformer-based models in this category utilize sequence-to-sequence frameworks to generate missing entities or relations in a knowledge graph directly.
-These models are pre-trained on large text corpora and fine-tuned on knowledge graph data, leveraging their ability to generate coherent and contextually appropriate text sequences.
-Each entity, relation, and triplet in the KG is treated as a text sequence. This allows for the use of textual prompts in LLMs for prediction tasks.
-$\text{Triple}=(v,e,w)$ (16)
+6 demonstrates the GenKGC [32] architecture that exemplifies structural encoding.
+Figure 6: An overview of structural encoding method.
+This method converts knowledge graph completion into a sequence-to-sequence generation task.
+It uses pre-trained language models to generate target entities from input sequences representing head entities and relations.
+The structural encoding is described as follows:
+Given a triple with a missing tail entity $(e_{h},r,?)$, the input sequence is constructed by concatenating the descriptions of the head entity $d_{e_{h}}$ and the relation $d_{r}$:
 […]
-For triple classification task, which involves determining if a given triple $(v,e,w)$ is correct (valid) or not, the prompt example is:
-[…]
-For triple classification task, Given $v$ and $w$, predict $e$,
-[…]
-For entity prediction task, depending on the missing entity, prompts are created to predict either $v$ or $w$, tbe prompt for predicting tail entity $w$ is:
-[…]
-The prompt for predicting head entity $v$ is:
-[…]
-Finally, uses instruction tuning with prompts and expected responses to fine-tune the models. This aligns the predictions with the factual correctness required for KG tasks.
-KGT5 52 is pre-trained and fine-tuned on Knowledge Graph Completion (KGC) and Knowledge Graph Question Answering (KGQA) tasks.
-The model employs regularization to maintain link prediction accuracy.
-The advantage of this method is its flexibility in handling both KGC and KGQA, making it a versatile solution.
-However, the complexity of training and fine-tuning a Seq2Seq model can be computationally expensive and time-consuming.
-KG-S2S 51 adopts a Seq2Seq generative framework to manage various types of KGC tasks by translating graph structures into flat text representations.
-This approach allows for unified processing but might suffer from a loss of structural information, which is critical for accurate graph completion.
-CSPromp-KG 32 leverages Conditional Soft Prompts to guide pre-trained language models in balancing the integration of structural and textual information for KGC.
-The method enhances model performance by ensuring a more nuanced understanding of graph structures.
-The major drawback is the potential overfitting to specific knowledge graph structures, which may limit the generalizability.
-The strength of Seq2Seq models lies in their flexibility and effectiveness in generating textual output that adheres to the structural rules of knowledge graphs.
-However, these models can be prone to generating implausible links if not properly constrained or if the training data is not representative enough of the test scenarios.
+where $d_{e_{h}}$ and $d_{r}$ are the textual descriptions of the head entity and the relation, respectively. The output sequence corresponds to the target tail entity $d_{e_{t}}$.
+The generative process is defined as:
+$p_{\theta}(y|x)=\prod_{i=1}^{N}p_{\theta}(y_{i}|y_{<i},x)$ (16)
+where $x$ is the input sequence, $y$ is the output sequence, and $N$ is the length of the output sequence. The model is trained using a standard sequence-to-sequence objective function:
+$L=-\log p_{\theta}(y|x)$ (17)
+During training, GenKGC employs relation-guided demonstration and entity-aware hierarchical decoding to improve representation learning and reduce inference time.
+Relation-guided demonstration augments the input sequence with examples of triples that share the same relation:
+$x=[\text{CLS}]\;\text{demonstration}(r)\;[\text{SEP}]\;d_{e_{h}}\;[\text{SEP}]\;d_{r}\;[\text{SEP}]$ (18)
+Entity-aware hierarchical decoding constrains the decoding process by utilizing entity type information, which helps in reducing the search space during generation and improving inference efficiency.
+LambdaKG [33] advances the application of PLM in knowledge graph embeddings.
+Unlike earlier methods that merely transform triples into embeddings, LambdaKG encapsulates both structural and textual information into a unified model architecture.
+It harnesses sophisticated training techniques like prompt engineering, negative sampling, and relation-aware modeling to enhance the efficiency and accuracy of knowledge graph representations.
+3.2.2 Fine-tuning
+Fine-tuning methods adapt pre-trained encoder-decoder models using natural language descriptions of triples.
+This approach tailors the model’s generative capabilities to specific knowledge representation tasks, improving performance on downstream applications.
+An illustration of a typical fine-tuning model KGT5 [34] is presented in Fig.
+7.
+Figure 7: An overview of fine-tuning method.
+KGT5 fine-tunes a T5 model on both link prediction and question answering tasks.
+The process involves first pre-training on link prediction and then fine-tuning on QA datasets, balancing between these tasks with a regularization approach.
+Given a set of entities $E$ and relations $R$, a knowledge graph $K\subseteq E\times R\times E$ consists of triples $(s,p,o)$.
+KGT5 treats KG link prediction and QA as sequence-to-sequence tasks.
+The model uses textual representations of entities and relations and a verbalization scheme to convert link prediction queries into textual queries.
+The training process can be described with the following formulas:
+$\text{Score}(a)=p_{a}+\alpha\quad\text{if}\;a\in N(e)$ (19)
+$\text{Score}(a)=p_{a}\quad\text{otherwise}$ (20)
+where $p_{a}$ is the log probability of the predicted entity $a$, $\alpha$ is a constant hyperparameter, and $N(e)$ represents the $n$-hop neighborhood of the entity $e$.
+The model’s architecture and training process can significantly reduce the model size while maintaining or improving performance on large KG and KGQA tasks.
+KG-S2S [35] is a state-of-the-art Seq2Seq generative framework designed to tackle various challenges in knowledge graph completion (KGC) across different settings without necessitating structural modifications for each graph type.
+This model uniquely addresses the limitation of previous KGC approaches that are tightly coupled with specific graph structures, which restricts their adaptability to new or evolving KG.
+By treating all elements of knowledge graphs—entities, relations, and even metadata—as sequences in a unified ‘flat’ text format, KG-S2S avoids the complexity inherent in traditional graph-based methods.
+This innovation not only simplifies the input data representation but also enhances the model’s flexibility and scalability.
+It employs advanced fine-tuning techniques on pre-trained language models, integrating novel mechanisms such as entity descriptions and relational soft prompts to enrich the model’s contextual understanding.
 3.3 Decoder-Based Methods
-In these studies, LLMs are adapted to enhance reasoning and inference capabilities over knowledge graphs, particularly for complex question answering and fact verification tasks.
-Techniques include adapting attention mechanisms to focus on relevant subgraphs and employing reasoning strategies that leverage structured graph data.
-3.3.1 Multi-Level Encoder
-Fig. LABEL:fig2.2 illustrates the workflow of this approach.
-This category includes models that enhance the capabilities of LLMs in complex reasoning and question answering tasks over knowledge graphs.
-Techniques involve structuring the input to LLMs in a way that mimics reasoning or using retrieval-augmented generation to enhance the context available for decision-making.
-The approach starts by converting a relevant subgraph $\mathcal{G}_{q}$ from the knowledge graph into a sequence using a breadth-first search (BFS)-based serialization.
-This serialized form is then processed by the PLM.
-The serialized subgraph $\mathcal{G}_{q}$ is denoted as:
-$\mathcal{SG}_{q}=\{v,e,w,\ldots\}$ (21)
-where $v$, $e$, and $w$ are nodes and edges in the subgraph.
-As for subgraph-aware self-attention, the mechanism is designed to mimic the processing of a GNN, enabling the model to understand the structure of the subgraph and reason over it.
-It utilizes attention mechanisms to model the relations and interactions between different entities (nodes) and relations (edges) within the subgraph.
-The attention process can be represented as:
-$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V$ (22)
-Here, $Q$, $K$, and $V$ are query, key, and value matrices, respectively, and $M$ is a mask matrix that directs attention flow to maintain structural integrity.
-The PLM is adapted to handle the serialized subgraph and question data simultaneously, facilitating deep interaction and knowledge sharing between the question context and the graph structure.
-This integration enables the PLM to perform reasoning tasks directly on the subgraph structure, harnessing both the rich semantic understanding of the PLM and the structured reasoning capabilities typically associated with GNNs.
-The PLM undergoes adaptation tuning using a dataset created from synthesized questions and subgraphs.
-This adaptation helps the PLM better understand and process the unique format of KG data.
-The fine-tuning phase optimizes the PLM on downstream question answering tasks over KGs, improving accuracy and efficiency in real-world applications.
-ReSKGC 81 introduces a retrieval-enhanced Seq2Seq model that uses a retrieval module to enhance the generation of missing entities by fetching semantically relevant triplets.
-This method improves accuracy and relevance but relies heavily on the effectiveness of the retrieval process, which can be a bottleneck if not optimized properly.
-StructGPT 88 employs an Iterative Reading-then-Reasoning framework to enhance LLMs’ reasoning capabilities over structured data like knowledge graphs.
-It significantly improves reasoning performance, but the iterative process can be resource-intensive and slow, particularly with complex queries.
-These models excel in tasks requiring deep reasoning and can handle nuanced queries effectively by leveraging both the structured information from knowledge graphs and the unstructured textual data processed by LLMs.
-The main drawback is their potential computational inefficiency and the complexity of integrating structured reasoning with generative language models.
+Decoder-based methods, utilizing models such as LLaMA [36] and GPT-4 [37], are characterized by the role of the decoder in the representation learning process.
+Description Generate methods (§3.3.1) create descriptive text for low-resource entities while using Encoder-based or Encoder-Decoder-based methods for other tasks.
+Prompt Engineering (§3.3.2)treats the decoder as a question-answering tool, using natural language to retrieve triple information and perform downstream tasks.
+Structural Fine-tuning (§3.3.3) combines structural and textual embeddings, optimizing the decoder’s output for improved knowledge representation.
+3.3.1 Description Generate
+Description generate methods enhance the representation of low-resource entities by creating descriptive text.
+This approach supplements the textual information, improving the performance of previous methods and ensuring a richer, more complete entity representation.
+Fig.
+8 illustrates a typical model CD [38] used in description generate method.
+Figure 8: An overview of description generate method.
+This approach leverages LLM to transform compact structural triplets into context-rich segments.
+The process involves generating descriptive contexts for knowledge graph completion using prompts.
+The key steps can be summarized as follows:
+Given a triplet $(h,r,t)$ where $h$ is the head entity, $r$ is the relation, and $t$ is the tail entity, the method generates descriptive context $c$ using an LLM prompted with the triplet.
+The process can be expressed as:
+$p_{i}=\text{Template}(h_{i},r_{i},t_{i})$ (21)
+$c_{i}=\text{LLM}(p_{i})$ (22)
+where $p_{i}$ is the prompt template filled with the triplet components, and $c_{i}$ is the descriptive context generated by the LLM.
+Two auxiliary tasks, reconstruction and contextualization, are introduced to train smaller KGC models with these enriched triplets.
+The reconstruction task aims to restore corrupted descriptive contexts using masked language modeling (MLM), while the contextualization task trains the model to generate descriptive context from the original triplet.
+The losses for these tasks are defined as:
+$L_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f(\text{MLM}(c_{i})),c_{i})$ (23)
+$L_{\text{con}}=\frac{1}{N}\sum_{i=1}^{N}\ell(f(I_{i}),c_{i})$ (24)
+where $\ell$ is the cross-entropy loss function, $f$ is the model, $I_{i}$ is the concatenated input of the head, relation, and tail, and $N$ is the number of samples.
+The final loss for training the KGC models combines the KGC loss with the auxiliary task losses:
+$L_{\text{final}}=L_{\text{kgc}}+\alpha\cdot L_{\text{rec}}+\beta\cdot L_{\text{con}}$ (25)
+where $\alpha$ and $\beta$ are hyperparameters that balance the contributions of the auxiliary losses.
+The CP-KGC [39] model represents a significant advancement in the realm of text-based KGC.
+It specifically leverages Constrained-Prompt Knowledge Graph Completion with Large Language Models to refine and enhance textual descriptions within KGC datasets.
+CP-KGC utilizes simple, carefully designed prompts that help in regenerating or supplementing existing textual descriptions, thereby improving the overall expressiveness and utility of the data.
+This approach effectively addresses issues like hallucinations in text generation by LLM, ensuring more accurate and contextually relevant outputs.
 3.3.2 Prompt Engineering
-Research in this area focuses on integrating comprehensive contextual information and structured knowledge into the training and operation of LLMs.
-This enables the models to better understand and generate responses based on both the textual descriptions and the underlying graph structures of the knowledge graph.
-The conventional method flowchart can be seen in Fig.
+Prompt engineering leverages the natural language capabilities of decoders, framing knowledge retrieval and representation tasks as question-answering problems.
+This method utilizes the decoder to extract and learn from knowledge graph triples efficiently.
+A representative model KG-LLM [40] using prompt engineering is shown in Fig.
 9.
 Figure 9: An overview of prompt engineering method.
-Methods under this heading focus on integrating extensive contextual information and structured knowledge graph data within the processing capabilities of LLMs.
-This often involves using knowledge prompts or other forms of guided input to enhance the LLM’s understanding of the graph structure and semantics.
-Complex questions are decomposed into simpler sub-questions based on predefined templates. This is achieved through a slot-filling mechanism driven by LLMs, formalized as:
-$\{\mathcal{S}_{i,t}\}_{t=1}^{T}=\mathrm{LLM}(\mathcal{Q}_{i})$ (23)
-where $\mathcal{Q}_{i}$ is the initial complex question, and $\mathcal{S}_{i,t}$ represents the decomposed sub-questions.
-For each decomposed sub-question, the framework retrieves relevant entities and facts from the KG using predefined logical chains that correspond to the sub-questions, the retrieval process is summarized as:
-$\mathcal{C}_{t}=\bigcup_{l=1}^{L_{i}}\{(s,r,o)_{l}\mid(s,r,o)_{l}\in\mathcal{G}\}$ (24)
-where $(s,r,o)_{l}$ are the triples retrieved based on the logical chains, with $s$, $r$, and $o$ denoting the subject, relation, and object, respectively.
-The candidate reasoning stage involves selecting the correct answers from the retrieved candidates by evaluating their relevance to the sub-questions, expressed using the following function:
-$\mathcal{A}_{i}=\mathrm{LLM}(\mathcal{Q}_{i,t}\mid\mathcal{C}_{t})$ (25)
-where $\mathcal{A}_{i}$ are the answers derived for sub-question $\mathcal{Q}_{i,t}$, based on candidates $\mathcal{C}_{t}$.
-Finally, the responses from all sub-questions are aggregated to construct a comprehensive answer to the original question, which formalized as:
-$\text{Response}=\text{Generate}(\{\mathcal{A}_{i}\}_{i=1}^{N})$ (26)
-where the function Generate compiles the individual answers into a final coherent response.
-KoPA 56 introduces the Knowledge Prefix Adapter (KoPA), which enhances the integration of structural KG information into LLMs for better structural-aware reasoning.
-While it improves reasoning accuracy, the adaptation of LLMs to specific knowledge graph structures could limit the model versatility across different KGs.
-CP-KGC 31 explores the use of zero-shot large language models to enhance text-based KGC without fine-tuning.
-The advantage is the minimal training requirement, but the reliance on zero-shot capabilities may result in less precise or contextually inappropriate text generation.
-KICGPT 53 integrates a triple-based knowledge graph completion retriever with a large language model to address information scarcity in long-tail entities.
-This method effectively utilizes in-context learning strategies, but its success heavily depends on the quality and relevance of the retrieved knowledge.
-These approaches allow LLMs to leverage their pre-trained knowledge effectively by guiding them with structured inputs, enhancing their performance on knowledge-intensive tasks.
-However, the success of these methods heavily depends on the quality and the extent of the integration between text and graph data, which can be challenging to optimize.
+This method transforms triples into natural language prompts and uses LLM to predict the plausibility of the triples or complete missing information. The process can be described as follows:
+For a given triple $(h,r,t)$, where $h$ is the head entity, $r$ is the relation, and $t$ is the tail entity, the prompt is constructed to query the LLM.
+For instance, for triple classification, the prompt might be:
+[…]
+The LLM then generates a response that indicates the plausibility of the triple. The model uses a scoring function to quantify the plausibility of the triple:
+$s(h,r,t)=\text{LLM}(p(h,r,t))$ (27)
+where $p(h,r,t)$ is the prompt generated for the triple $(h,r,t)$, and LLM is the large language model.
+The training objective includes a cross-entropy loss to fine-tune the LLM on labeled triples:
+$L=-\sum_{(h,r,t)\in D}[y\log s(h,r,t)+(1-y)\log(1-s(h,r,t))]$ (28)
+where $D$ is the set of training triples, and $y$ is the label indicating whether the triple is true or false.
+The KICGPT [41] model integrates a LLM with KGC methods to address challenges in traditional KGC approaches.
+It significantly reduces training overhead by utilizing in-context learning strategies without requiring explicit model fine-tuning.
+The model leverages the LLM’s extensive pre-trained knowledge base, combined with a structure-aware KG retriever, to improve the handling of long-tail entities.
+This integration allows the KICGPT to effectively utilize both structured KG information and the broad knowledge base of the LLM, providing a more robust framework for KGC tasks.
+The model is particularly innovative in its use of ‘Knowledge Prompt’ strategies, which guide the LLM through structured prompts that include KG information, enhancing the model’s ability to make informed predictions about missing entities in KG.
+3.3.3 Structural Fine-tuning
+Structural fine-tuning combines embeddings from both structural and textual information, feeding them into the decoder.
+This approach aims to optimize the decoder’s output by providing a more comprehensive and nuanced representation of the knowledge graph triples.
+Fig.
+10 depicts a structural fine-tuning model KoPA [42].
+Figure 10: An overview of structural fine-tuning method.
+This method integrates structural embeddings of entities and relations into LLM to enhance knowledge graph completion tasks.
+The core idea of KoPA involves two main steps: pre-training structural embeddings and using a prefix adapter to inject these embeddings into the LLM.
+The structural embeddings are first learned through a self-supervised pre-training process, capturing the structural information of the knowledge graph.
+The process is described as follows:
+For each triple $(h,r,t)$, the structural embeddings $\mathbf{h},\mathbf{r},\mathbf{t}$ are learned using the scoring function:
+$F(h,r,t)=\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|^{2}$ (29)
+The pre-training objective involves a margin-based ranking loss with negative sampling:
+$L_{\text{pre}}=-\log\sigma(\gamma-F(h,r,t))-\sum_{i=1}^{K}\log\sigma(F(h^{\prime}_{i},r^{\prime}_{i},t^{\prime}_{i})-\gamma)$ (30)
+where $(h^{\prime}_{i},r^{\prime}_{i},t^{\prime}_{i})$ are negative samples, $\gamma$ is the margin, and $\sigma$ is the sigmoid function.
+Once the structural embeddings are pre-trained, KoPA employs a prefix adapter to map these embeddings into the textual token space of the LLM. The transformation is performed as follows:
+$\mathbf{K}=P(\mathbf{h})\oplus P(\mathbf{r})\oplus P(\mathbf{t})$ (31)
+where $P$ is a projection layer, and $\oplus$ denotes concatenation. The transformed embeddings, referred to as virtual knowledge tokens, are prepended to the input sequence $S$:
+$S_{\text{KoPA}}=\mathbf{K}\oplus I\oplus X$ (32)
+where $I$ is the instruction prompt, and $X$ is the triple prompt.
+The training objective for fine-tuning the LLM with KoPA involves minimizing the cross-entropy loss:
+$L_{\text{KoPA}}=-\frac{1}{|S_{\text{KoPA}}|}\sum_{i=1}^{|S_{\text{KoPA}}|}\log P_{\text{LLM}}(s_{i}|s_{<i})$ (33)
+where $s_{i}$ are the tokens in the input sequence $S_{\text{KoPA}}$.
+KG-GPT2 [43] is a novel adaptation of the GPT-2 language model for KGC.
+KG-GPT2 leverages the contextual capabilities of GPT-2 to address the challenge of incomplete knowledge graphs by predicting missing links and relationships within them.
+This model treats each triple in a knowledge graph as a sentence, allowing the GPT-2’s language understanding capabilities to classify the likelihood of these triples accurately.
+By contextualizing triples in this manner, KG-GPT2 goes beyond traditional embedding techniques, incorporating richer linguistic and semantic information to enhance the graph’s completion.
 4 Experiments and Evaluations
-In this section, we present a comprehensive suite of experiments and evaluations to assess the performance and effectiveness of the various LLM-enhanced KRL methods discussed previously.
+In this section, we present a comprehensive suite of experiments and evaluations to assess the performance and effectiveness of the various LLM enhanced KRL methods discussed previously.
 We utilize a variety of datasets and metrics to provide a thorough analysis, ensuring that our findings are robust and generalizable.
@@
 Table 4: Statistics on Datasets.
-Dataset # Entity # Relation # Train # Dev # Test Task Year Source FIGER - - 2,000,000 10,000 563 ET 2015 1 Open Entity - - 2,000 2,000 2,000 ET 2018 2 FewRel 1.0 - 80 8,000 16,000 16,000 RC 2018 3 TACRED - 42 68,124 22,631 15,509 RC 2017 4 WN11 38,696 11 112,581 2,609 10,544 TC 2013 5 WN9 6,555 9 11,741 1,337 1,319 TC 2022 6 FB13 75,043 13 316,232 5,908 23,733 TC 2013 5 FB15K-237N 13,104 93 87,282 7,041 8,226 TC 2022 7 FB15K 14,951 1,345 483,142 50,000 59,071 LP 2013 8 FB15k-237 14,541 237 272,115 17,535 20,466 LP 2015 9 UMLS 135 46 5,216 652 661 LP 2018 10 Nations 14 55 1,592 199 201 LP 2007 11 Diabetes 7,886 67 56,830 1,344 1,936 LP 2022 2 YAGO3-10 103,222 30 490,214 2,295 2,292 LP 2018 10 Wikidata5M (transductive) 4,594,485 822 20,614,279 5,163 5,133 LP 2021 13 Wikidata5M (inductive) 4,594,458 822 20,496,514 6,699 6,894 LP(ZS) CoDEx-S 45,869 68 32,888 1,827 1,828 LP 2020 14 CoDEx-M 11,941 50 185,584 10,310 10,311 LP CoDEx-L 45,869 69 551,193 30,622 30,622 LP NELL-ONE 68,545 822 189,635 1,004 2,158 LP(ZS) 2018 15 WN18RR 40,943 11 86,835 3,034 3,134 LP, LP(ZS) 2018 10
+Dataset # Ent # Rel # Train # Dev # Test Task Year
+FIGER [44] - - 2,000,000 10,000 563 ET 2015
+Open Entity [45] - - 2,000 2,000 2,000 ET 2018
+TACRED [46] - 42 68,124 22,631 15,509 RC 2017
+FewRel [47] - 80 8,000 16,000 16,000 RC 2018
+WN11 [48] 38,696 11 112,581 2,609 10,544 TC 2013
+FB13 [48] 75,043 13 316,232 5,908 23,733 TC 2013
+WN9 [49] 6,555 9 11,741 1,337 1,319 TC 2022
+FB15K-237N [50] 13,104 93 87,282 7,041 8,226 TC 2022
+Nations [51] 14 55 1,592 199 201 LP 2007
+FB15K [52] 14,951 1,345 483,142 50,000 59,071 LP 2013
+FB15k-237 [53] 14,541 237 272,115 17,535 20,466 LP 2015
+WN18RR [54] 40,943 11 86,835 3,034 3,134 LP, LP(ZS) 2018
+UMLS [54] 135 46 5,216 652 661 LP 2018
+YAGO3-10 [54] 103,222 30 490,214 2,295 2,292 LP 2018
+NELL-ONE [55] 68,545 822 189,635 1,004 2,158 LP(ZS) 2018
+CoDEx-S [56] 45,869 68 32,888 1,827 1,828 LP 2020
+CoDEx-M [56] 11,941 50 185,584 10,310 10,311 LP 2020
+CoDEx-L [56] 45,869 69 551,193 30,622 30,622 LP 2020
+Wikidata5M-transductive [28] 4,594,485 822 20,614,279 5,163 5,133 LP 2021
+Wikidata5M-inductive [28] 4,594,458 822 20,496,514 6,699 6,894 LP(ZS) 2021
+Diabetes [57] 7,886 67 56,830 1,344 1,936 LP 2022
   • FIGER.
@@
 It provides annotations for entities within free text, helping to improve the classification of entities into a broad set of predefined categories, enhancing the ability to understand and process natural language.
+  • TACRED.
+TACRED (TAC Relation Extraction Dataset) is a large-scale relation extraction dataset created by Stanford.
+It includes sentences annotated with a wide range of relation types, providing a rich resource for training and evaluating models on the task of extracting relations between entities in text.
   • FewRel.
@@
 It includes various relation types and supports the development of models that can learn to classify relations from a few instances.
-  • TACRED.
-TACRED (TAC Relation Extraction Dataset) is a large-scale relation extraction dataset created by Stanford.
-It includes sentences annotated with a wide range of relation types, providing a rich resource for training and evaluating models on the task of extracting relations between entities in text.
   • WN11.
@@
 It includes 13 relation types and provides a benchmark for evaluating the ability of models to predict missing links in a knowledge graph.
+  • WN9.
+WN9 is another subset of the WordNet knowledge graph, designed for tasks such as knowledge graph completion and link prediction.
+It includes a specific set of entities and relations to facilitate focused research in these areas.
+  • Nations.
+The Nations dataset is a small-scale dataset used for relation prediction and knowledge graph completion.
+It includes entities representing countries and various relations among them, offering a simplified benchmark for evaluating models.
   • FB15K.
@@
 It excludes inverse relations, making it a more challenging and realistic benchmark for knowledge graph completion tasks.
-  • WN18.
-WN18 is a benchmark dataset based on the WordNet knowledge graph.
-It contains a subset of entities and relations, often used for evaluating models on the task of link prediction in knowledge graphs.
-  • WN9.
-WN9 is another subset of the WordNet knowledge graph, designed for tasks such as knowledge graph completion and link prediction.
-It includes a specific set of entities and relations to facilitate focused research in these areas.
   • WN18RR.
@@
 It supports research in few-shot relation learning by providing a challenging benchmark with a variety of relation types annotated from web text.
+  • CoDEx.
+CoDEx (Comprehensive Knowledge Graph Completion and Explanation) is a suite of knowledge graph completion datasets designed to provide diverse and challenging benchmarks.
+It includes multiple subsets with different relation types and complexities, supporting research in explainable AI and knowledge graph reasoning.
   • Wikidata5M.
@@
 It is used for various tasks, including knowledge graph completion, entity linking, and question answering, providing a comprehensive resource for natural language understanding.
-  • Nations.
-The Nations dataset is a small-scale dataset used for relation prediction and knowledge graph completion.
-It includes entities representing countries and various relations among them, offering a simplified benchmark for evaluating models.
-  • CoDEx.
-CoDEx (Comprehensive Knowledge Graph Completion and Explanation) is a suite of knowledge graph completion datasets designed to provide diverse and challenging benchmarks.
-It includes multiple subsets with different relation types and complexities, supporting research in explainable AI and knowledge graph reasoning.
 4.2 Metrics
@@
 Table 5: Experimental statistics under the ET subtask.
-Model Year FIGER Open Entity Choi et al.
-Acc Macro Micro P R F1 P R F1 NFGEC (Attentive) 70 2016 54.53 74.76 71.58 - - - - - - NFGEC (LSTM) 70 2016 55.6 75.15 71.73 68.8 53.3 60.1 - - - UFET 98 2018 - - - 77.4 60.6 68 68.8 53.3 60.1 BERT 18 2018 52.04 75.16 71.63 76.37 70.96 73.56 76.4 71 73.6 RoBERTa 83 2019 - - - 77.4 73.6 75.4 - - - ERNIE 38 2019 57.19 76.51 73.39 78.42 72.9 75.56 78.4 72.9 75.6 KnowBert 55 2019 - - - 78.7 72.7 75.6 78.6 73.7 76.1 KEPLER 47 2021 - - - 77.8 74.6 76.2 - - -
+Model Year FIGER Open Entity
+Acc Macro Micro P R F1
+NFGEC [58] 2016 55.6 75.15 71.73 68.8 53.3 60.1
+UFET [45] 2018 - - - 77.4 60.6 68
+BERT [1] 2018 52.04 75.16 71.63 76.37 70.96 73.56
+RoBERTa [6] 2019 - - - 77.4 73.6 75.4
+ERNIE [59] 2019 57.19 76.51 73.39 78.42 72.9 75.56
+KnowBert [60] 2019 - - - 78.7 72.7 75.6
+KEPLER [28] 2021 - - - 77.8 74.6 76.2
 The transformer-based models’ superior performance highlights their effectiveness in ET tasks.
@@
 Table 6: Experimental statistics under the RC subtask.
-Model Year FewRel 1.0 TACRED P R F1 P R F1 CNN 23 2015 69.51 69.64 69.35 70.3 54.2 61.2 PA-LSTM 73 2017 - - - 65.7 64.5 65.1 C-GCN 21 2018 - - - 69.9 63.3 66.4 BERT_Base 18 2018 85.05 85.11 84.89 67.23 64.81 66 BERT_Large 18 2018 - - - - - 70.1 RoBERTa 83 2019 - - - 70.4 71.1 70.7 MTB 65 2019 - - - - - 71.5 ERNIE 38 2019 88.49 88.44 88.32 69.97 66.08 67.97 KnowBert 55 2019 - - - 71.6 71.4 71.5 KEPLER 47 2021 - - - 71.5 72.5 72
+Model Year FewRel TACRED
+P R F1 P R F1
+CNN [61] 2015 69.51 69.64 69.35 70.3 54.2 61.2
+PA-LSTM [62] 2017 - - - 65.7 64.5 65.1
+C-GCN [63] 2018 - - - 69.9 63.3 66.4
+BERT_Base [1] 2018 85.05 85.11 84.89 67.23 64.81 66
+BERT_Large [1] 2018 - - - - - 70.1
+RoBERTa [6] 2019 - - - 70.4 71.1 70.7
+MTB [64] 2019 - - - - - 71.5
+ERNIE [59] 2019 88.49 88.44 88.32 69.97 66.08 67.97
+KnowBert [60] 2019 - - - 71.6 71.4 71.5
+KEPLER [28] 2021 - - - 71.5 72.5 72
 The transformer-based models’ superior performance highlights their effectiveness in RC tasks.
@@
 Table 7: Experimental statistics under the RP subtask.
-Model Year FB15K YAGO3-10-100 MR Hits@1 Hits@1 TransE 92 2013 2.5 84.3 - TransR 96 2015 2.1 91.6 - DKRL 36 2016 2 90.8 - TKRL 90 2016 1.7 92.8 - PTransE 76 2015 1.2 93.6 - SSP 87 2017 1.2 - - ProjE 75 2017 1.2 95.7 - KG-BERT 48 2019 1.2 96 - KGT5 52 2022 - - 60 ChatGPT 2023 - - 39 GPT-4 2023 - - 56 LLaMA-7B 59 2023 - - 13 LLaMA-13B 59 2023 - - 1 KG-LLM 50 2023 - - 71
+Model Year FB15K YAGO3-10-100
+MR Hits@1 Hits@1
+TransE [52] 2013 2.5 84.3 -
+TransR [65] 2015 2.1 91.6 -
+DKRL [66] 2016 2 90.8 -
+TKRL [67] 2016 1.7 92.8 -
+PTransE [68] 2015 1.2 93.6 -
+SSP [69] 2017 1.2 - -
+ProjE [70] 2017 1.2 95.7 -
+KG-BERT [19] 2019 1.2 96 -
+KGT5 [34] 2022 - - 60
+ChatGPT 2023 - - 39
+GPT-4 2023 - - 56
+LLaMA-7B [36] 2023 - - 13
+LLaMA-13B [36] 2023 - - 1
+KG-LLM [40] 2023 - - 71
 Transformer-based models like KG-BERT and KG-GPT2 leverage the power of pre-trained language models, which can capture intricate semantic relationships and contextual nuances within triples.
@@
 Table 8: Experimental statistics under the TC subtask.
-Model Year WN11 FB13 UMLS CoDEX-S FB15k-237N Acc Acc Acc Acc Acc NTN 71 2013 86.2 90 - - - TransE 92 2013 75.9 81.5 78.1 72.1 - DistMult 34 2014 87.1 86.2 86.8 66.8 - ComplEx 25 2016 - - 90.8 67.8 65.7 RotatE 84 2019 - - 92.1 75.7 68.5 KG-BERT 48 2019 93.5 90.4 89.7 77.3 56 KG-GPT2 49 2021 85 89 - - - LaSS 57 2021 94.5 91.8 - - - KGT5 52 2022 72.8 66.3 - - - LMKE 60 2022 - 91.7 92.4 - - GPT-3.5 2023 - - 67.6 54.7 60.2 LLaMA-7B 59 2023 21.1 9.1 - - - LLaMA-13B 59 2023 28.1 17.6 - - - KG-LLM 50 2023 95.6 90.2 86 80.3 80.5 KoPA 56 2023 - - 92.6 82.7 77.7
+Model Year WN11 FB13 UMLS CoDEX-S FB15k-237N
+Acc Acc Acc Acc Acc
+NTN [48] 2013 86.2 90 - - -
+TransE [52] 2013 75.9 81.5 78.1 72.1 -
+DistMult [71] 2014 87.1 86.2 86.8 66.8 -
+ComplEx [72] 2016 - - 90.8 67.8 65.7
+RotatE [73] 2019 - - 92.1 75.7 68.5
+KG-BERT [19] 2019 93.5 90.4 89.7 77.3 56
+KG-GPT2 [43] 2021 85 89 - - -
+LaSS [74] 2021 94.5 91.8 - - -
+KGT5 [34] 2022 72.8 66.3 - - -
+LMKE [75] 2022 - 91.7 92.4 - -
+GPT-3.5 2023 - - 67.6 54.7 60.2
+LLaMA-7B [36] 2023 21.1 9.1 - - -
+LLaMA-13B [36] 2023 28.1 17.6 - - -
+KG-LLM [40] 2023 95.6 90.2 86 80.3 80.5
+KoPA [42] 2023 - - 92.6 82.7 77.7
 The performance of these models highlights several trade-offs and considerations:
@@
 This task is essential for expanding and enriching knowledge graphs, making them more useful for various applications like recommendation systems and semantic search.
-Table 9: Experimental statistics of WN18RR and FB15k-237 datasets under the LP subtask in transductive settings.
-Model Year WN18RR FB15k-237 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 2300 24.3 4.3 44.1 53.2 223 27.9 19.8 37.6 47.4 TransH 94 2014 2524 - - - 50.3 255 - - - 48.6 DistMult 34 2014 3704 44.4 41.2 47 50.4 411 28.1 199 30.1 44.6 TransR 96 2015 3166 - - - 50.7 237 - - - 51.1 TransD 91 2015 2768 - - - 50.7 246 - - - 48.4 ComplEx 25 2016 3921 44.9 40.9 46.9 53 508 27.8 19.4 29.7 45 ConvE 29 2018 4464 45.6 41.9 47 53.1 245 31.2 22.5 34.1 49.7 ConvKB 30 2018 2554 24.9 - - 52.5 257 24.3 - - 51.7 R-GCN 82 2018 6700 12.3 8 13.7 20.7 600 16,4 10 18.1 41.7 KBGAN 46 2018 - 21.5 - - 48.1 - 27.7 - - 45.8 RotatE 84 2019 3340 47.6 42.8 49.2 57.1 177 33.8 24.1 37.5 53.3 KBAT 45 2019 1921 41.2 - - 55.4 270 15.7 - - 33.1 CapsE 19 2019 718 41.5 - - 55.9 403 15 - - 35.6 QuatE 78 2019 3472 48.1 43.6 50 56.4 176 31.1 22.1 34.2 49.5 TuckER 97 2019 - 47 44.3 48.2 52.6 - 35.8 26.6 39.4 54.4 HAKE 44 2019 - 49,7 45.2 51.6 58.2 - 34.6 25 38.1 54.2 AttH 17 2020 - 48.6 44.3 49.9 57.3 - 34.8 25.2 38.4 54 REFE 79 2020 - - - - 56.1 - - - - 54.1 GAATs 39 2020 1270 - - - 60.4 187 - - - 65 Complex-DURA 26 2020 57.1 - - - - 56 - - - - DensE 33 2020 3052 49.1 44.3 50.8 57.9 169 34.9 25.6 38.4 53.5 LineaRE 58 2020 1644 49.5 45.3 50.9 57.8 155 35.7 26.4 39.1 54.5 RESCAL-DURA 80 2020 - 49.8 45.5 - 57.7 - 36.8 27.6 - 55 CompGCN 24 2020 - 47.9 44.3 49.4 54.6 - 35.5 26.4 39 53.5 NePTuNe 69 2021 - - - - 55.7 - - - - 54.7 ComplEx-N3-RP 27 2021 - - - - 58 - - - - 56.8 ConE 28 2021 - 49.6 45.3 51.5 57.9 - 34.5 24.7 38.1 54 Rot-Pro 85 2021 - 45.7 39.7 48.2 57.7 - 34.4 24.6 38.3 54 QuatDE 77 2021 1977 48.9 43.8 50.9 58.6 90 36.5 26.8 40 56.3 NBFNet 68 2021 - 55.1 49.7 - 66.6 - 41.5 32.1 - 59.9 KG-BERT 48 2019 97 21.6 4.1 30.2 52.4 153 23.7 16.9 26 42.7 MTL-KGC 66 2020 89 33.1 20.3 38.3 59.7 132 26.7 17.2 29.8 45.8 Pretrain-KGE 74 2020 - 48.8 43.7 50.9 58.6 - 35 25 38.4 55.4 StAR 88 2021 51 40.1 24.3 49.1 70.9 1117 29.6 20.5 32.2 48.2 MEM-KGC 63 2021 - 57.2 48.9 62 72.3 - 34.9 26 38.2 52.4 LaSS 57 2021 35 - - - 78.6 108 - - - 53.3 SimKGC 86 2021 - 66.7 58.8 72.1 80.5 - 33.6 24.9 36.2 51.1 LP-BERT 62 2022 92 48.2 34.3 56.3 75.2 154 31 22.3 33.6 49 Nayyeri et al.
-67 2022 - - - - - - 33.6 24.5 37 52.2 KGT5 52 2022 - 54.2 50.7 - 60.7 - 34.3 25.2 - 37.7 OpenWorld KGC 72 2022 - 55.7 47.5 60.4 70.4 - 34.6 25.3 38.1 53.1 Lovelace et al.
-61 2022 - 59.1 51.8 61.6 73.5 - 37.5 28.3 40.8 56.4 LMKE 60 2022 79 61.9 52.3 67.1 78.9 141 30.6 21.8 33.1 48.4 GenKGC 40 2022 - - 28.7 40.3 53.5 - - 19.2 35.5 43.9 KG-S2S 51 2022 - 57.4 53.1 59.5 66.1 - 33.6 25.7 37.3 49.8 kNN-KGE 54 2023 - 57.9 52.5 - - - 28 37.3 - - CSPromp-KG 32 2023 - 57.5 52.2 59.6 67.8 - 35.8 26.9 39.3 53.8 GPT-3.5 2023 - - 19 - - - - 23.7 - - LLaMA-7B 59 2023 - - 8.5 - - - - - - - LLaMA-13B 59 2023 - - 10 - - - - - - - KG-LLM 50 2023 - - 25.59 - - - - - - - CD 20 2024 - 57.6 52.6 60.7 67.2 - - - - - CP-KGC 31 2024 - 67.3 59.9 72.1 80.4 - 33.8 25.1 36.5 51.6 KICGPT 53 2024 - 56.4 47.8 61.2 67.7 - 41.2 32.7 44.8 55.4
-Table 10: Experimental statistics of UMLS, Wikidata5M, and FB15k-237N datasets under the LP subtask in transductive setting.
-Model Year UMLS Wikidata5M FB15k-237N MR Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 1.84 98.9 109370 25.3 17 31.1 39.2 25.5 15.2 30.1 45.9 TransH 94 2014 1.8 99.5 - - - - - - - - - DistMult 34 2014 5.52 84.6 - 25.3 20.9 27.8 33.4 20.9 14.3 23.4 33 TransR 96 2015 1.81 99.4 - - - - - - - - - TransD 91 2015 1.71 99.3 - - - - - - - - - ComplEx 25 2016 2.59 96.7 - 30.8 25.5 - 39.8 24.9 18 27.6 38 DKRL 36 2016 - - 31566 16 12 18.1 22.9 - - - - ConvE 29 2018 1.51 99 - - - - - 27.3 19.2 30.5 42.9 RoBERTa 83 2019 - - 1381597 0.1 0 0.1 0.3 - - - - RotatE 84 2019 - - - 29 23.4 32.2 39 27.9 17.7 32 48.1 QuatE 78 2019 - - - 27.6 22.7 30.1 35.9 - - - - CompGCN 24 2020 - - - - - - - 31.6 23.1 34.9 48 NePTuNe 69 2021 - - - - - - - - - - - ComplEx-N3-RP 27 2021 - 99.8 - - - - - - - - - KG-BERT 48 2019 1.47 99 - - - - - 20.3 13.9 20.1 40.3 MTL-KGC 66 2020 - - - - - - - 24.1 16 28.4 43 StAR 88 2021 1.49 99.1 - - - - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - LaSS 57 2021 1.56 98.9 - - - - - - - - - KEPLER 47 2021 - - 200267 21 17.3 22.4 27.7 - - - - SimKGC 86 2021 - - - 35.8 31.3 37.6 44.1 - - - - LP-BERT 62 2022 1.18 100 - - - - - - - - - KGT5 52 2022 - - - 33.6 28.6 36.2 42.6 - - - - GenKGC 40 2022 - - - - - - - - 18.7 27.3 33.7 KG-S2S 51 2022 - - - - - - - 35.4 28.5 38.8 49.3 CSPromp-KG 32 2023 - - - 38 34.3 39.9 44.6 36 28.1 39.5 51.1 ReSKGC 81 2023 - - - 39.6 37.3 41.3 43.7 - - - - CD 20 2024 - - - - - - - 37.2 28.8 41 53
-Table 11: Experimental statistics under the LP subtask in inductive setting.
-Model Year NELL-One Wikidata5M N-Shot MRR Hits@1 Hits@5 Hits@10 N-Shot MRR MR Hits@1 Hits@3 Hits@10 DKRL 36 2016 - - - - - Zero-Shot 23.1 78 5.9 32 54.6 GMatching 41 2018 Five-Shot 20 14 26 31 - - - - - - MetaR 64 2019 Five-Shot 26 17 35 44 - - - - - - GMatching 41 2018 One-Shot 19 12 26 31 - - - - - - RoBERTa 83 2019 - - - - - Zero-Shot 7.4 723 0.7 1 19.6 MetaR 64 2019 One-Shot 25 17 34 40 - - - - - - StAR 88 2021 Zero-Shot 26 17 35 45 - - - - - - MEM-KGC 63 2021 - - - - - - - - - - - KEPLER 47 2021 - - - - - Zero-Shot 40.2 28 22.2 51.4 73 SimKGC 86 2021 - - - - - Zero-Shot 71.4 - 50.9 78.5 91.7 KG-S2S 51 2022 Zero-Shot 31 22 41 48 - - - - - -
+Table 9: Experimental statistics of WN18RR dataset under the LP subtask in transductive settings.
+Model Year WN18RR
+MR MRR Hits@1 Hits@3 Hits@10
+TransE [52] 2013 2300 24.3 4.3 44.1 53.2
+TransH [76] 2014 2524 - - - 50.3
+DistMult [71] 2014 3704 44.4 41.2 47 50.4
+TransR [65] 2015 3166 - - - 50.7
+TransD [77] 2015 2768 - - - 50.7
+ComplEx [72] 2016 3921 44.9 40.9 46.9 53
+ConvE [54] 2018 4464 45.6 41.9 47 53.1
+ConvKB [78] 2018 2554 24.9 - - 52.5
+R-GCN [79] 2018 6700 12.3 8 13.7 20.7
+KBGAN [80] 2018 - 21.5 - - 48.1
+RotatE [73] 2019 3340 47.6 42.8 49.2 57.1
+KBAT [81] 2019 1921 41.2 - - 55.4
+CapsE [82] 2019 718 41.5 - - 55.9
+QuatE [83] 2019 3472 48.1 43.6 50 56.4
+TuckER [84] 2019 - 47 44.3 48.2 52.6
+HAKE [85] 2019 - 49,7 45.2 51.6 58.2
+AttH [86] 2020 - 48.6 44.3 49.9 57.3
+REFE [87] 2020 - - - - 56.1
+GAATs [88] 2020 1270 - - - 60.4
+Complex-DURA [89] 2020 57.1 - - - -
+DensE [90] 2020 3052 49.1 44.3 50.8 57.9
+LineaRE [91] 2020 1644 49.5 45.3 50.9 57.8
+RESCAL-DURA [92] 2020 - 49.8 45.5 - 57.7
+CompGCN [93] 2020 - 47.9 44.3 49.4 54.6
+NePTuNe [94] 2021 - - - - 55.7
+ComplEx-N3-RP [95] 2021 - - - - 58
+ConE [96] 2021 - 49.6 45.3 51.5 57.9
+Rot-Pro [97] 2021 - 45.7 39.7 48.2 57.7
+QuatDE [98] 2021 1977 48.9 43.8 50.9 58.6
+NBFNet [99] 2021 - 55.1 49.7 - 66.6
+KG-BERT [19] 2019 97 21.6 4.1 30.2 52.4
+MTL-KGC [20] 2020 89 33.1 20.3 38.3 59.7
+Pretrain-KGE [100] 2020 - 48.8 43.7 50.9 58.6
+StAR [25] 2021 51 40.1 24.3 49.1 70.9
+MEM-KGC [101] 2021 - 57.2 48.9 62 72.3
+LaSS [74] 2021 35 - - - 78.6
+SimKGC [26] 2021 - 66.7 58.8 72.1 80.5
+LP-BERT [27] 2022 92 48.2 34.3 56.3 75.2
+KGT5 [34] 2022 - 54.2 50.7 - 60.7
+OpenWorld KGC [102] 2022 - 55.7 47.5 60.4 70.4
+LMKE [75] 2022 79 61.9 52.3 67.1 78.9
+GenKGC [32] 2022 - - 28.7 40.3 53.5
+KG-S2S [35] 2022 - 57.4 53.1 59.5 66.1
+kNN-KGE [103] 2023 - 57.9 52.5 - -
+CSPromp-KG [24] 2023 - 57.5 52.2 59.6 67.8
+GPT-3.5 2023 - - 19 - -
+LLaMA-7B [36] 2023 - - 8.5 - -
+LLaMA-13B [36] 2023 - - 10 - -
+KG-LLM [40] 2023 - - 25.59 - -
+CD [38] 2024 - 57.6 52.6 60.7 67.2
+CP-KGC [39] 2024 - 67.3 59.9 72.1 80.4
+KICGPT [41] 2024 - 56.4 47.8 61.2 67.7
+Table 10: Experimental statistics of FB15k-237 dataset under the LP subtask in transductive settings.
+Model FB15k-237
+MR MRR Hits@1 Hits@3 Hits@10
+TransE [52] 2013 223 27.9 19.8 37.6 47.4
+TransH [76] 2014 255 - - - 48.6
+DistMult [71] 2014 411 28.1 199 30.1 44.6
+TransR [65] 2015 237 - - - 51.1
+TransD [77] 2015 246 - - - 48.4
+ComplEx [72] 2016 508 27.8 19.4 29.7 45
+ConvE [54] 2018 245 31.2 22.5 34.1 49.7
+ConvKB [78] 2018 257 24.3 - - 51.7
+R-GCN [79] 2018 600 16,4 10 18.1 41.7
+KBGAN [80] 2018 - 27.7 - - 45.8
+RotatE [73] 2019 177 33.8 24.1 37.5 53.3
+KBAT [81] 2019 270 15.7 - - 33.1
+CapsE [82] 2019 403 15 - - 35.6
+QuatE [83] 2019 176 31.1 22.1 34.2 49.5
+TuckER [84] 2019 - 35.8 26.6 39.4 54.4
+HAKE [85] 2019 - 34.6 25 38.1 54.2
+AttH [86] 2020 - 34.8 25.2 38.4 54
+REFE [87] 2020 - - - - 54.1
+GAATs [88] 2020 187 - - - 65
+Complex-DURA [89] 2020 56 - - - -
+DensE [90] 2020 169 34.9 25.6 38.4 53.5
+LineaRE [91] 2020 155 35.7 26.4 39.1 54.5
+RESCAL-DURA [92] 2020 - 36.8 27.6 - 55
+CompGCN [93] 2020 - 35.5 26.4 39 53.5
+NePTuNe [94] 2021 - - - - 54.7
+ComplEx-N3-RP [95] 2021 - - - - 56.8
+ConE [96] 2021 - 34.5 24.7 38.1 54
+Rot-Pro [97] 2021 - 34.4 24.6 38.3 54
+QuatDE [98] 2021 90 36.5 26.8 40 56.3
+NBFNet [99] 2021 - 41.5 32.1 - 59.9
+KG-BERT [19] 2019 153 23.7 16.9 26 42.7
+MTL-KGC [20] 2020 132 26.7 17.2 29.8 45.8
+Pretrain-KGE [100] 2020 - 35 25 38.4 55.4
+StAR [25] 2021 1117 29.6 20.5 32.2 48.2
+MEM-KGC [101] 2021 - 34.9 26 38.2 52.4
+LaSS [74] 2021 108 - - - 53.3
+SimKGC [26] 2021 - 33.6 24.9 36.2 51.1
+LP-BERT [27] 2022 154 31 22.3 33.6 49
+KGT5 [34] 2022 - 34.3 25.2 - 37.7
+OpenWorld KGC [102] 2022 - 34.6 25.3 38.1 53.1
+LMKE [75] 2022 141 30.6 21.8 33.1 48.4
+GenKGC [32] 2022 - - 19.2 35.5 43.9
+KG-S2S [35] 2022 - 33.6 25.7 37.3 49.8
+kNN-KGE [103] 2023 - 28 37.3 - -
+CSPromp-KG [24] 2023 - 35.8 26.9 39.3 53.8
+GPT-3.5 2023 - - 23.7 - -
+CP-KGC [39] 2024 - 33.8 25.1 36.5 51.6
+KICGPT [41] 2024 - 41.2 32.7 44.8 55.4
+Table 11: Experimental statistics of UMLS and Wikidata5M datasets under the LP subtask in transductive setting.
+Model Year UMLS Wikidata5M
+MR Hits@10 MR MRR Hits@1 Hits@3 Hits@10
+TransE [52] 2013 1.84 98.9 109370 25.3 17 31.1 39.2
+TransH [76] 2014 1.8 99.5 - - - - -
+DistMult [71] 2014 5.52 84.6 - 25.3 20.9 27.8 33.4
+TransR [65] 2015 1.81 99.4 - - - - -
+TransD [77] 2015 1.71 99.3 - - - - -
+ComplEx [72] 2016 2.59 96.7 - 30.8 25.5 - 39.8
+DKRL [66] 2016 - - 31566 16 12 18.1 22.9
+ConvE [54] 2018 1.51 99 - - - - -
+RoBERTa [6] 2019 - - 1381597 0.1 0 0.1 0.3
+RotatE [73] 2019 - - - 29 23.4 32.2 39
+QuatE [83] 2019 - - - 27.6 22.7 30.1 35.9
+ComplEx-N3-RP [95] 2021 - 99.8 - - - - -
+KG-BERT [19] 2019 1.47 99 - - - - -
+StAR [25] 2021 1.49 99.1 - - - - -
+LaSS [74] 2021 1.56 98.9 - - - - -
+KEPLER [28] 2021 - - 200267 21 17.3 22.4 27.7
+SimKGC [26] 2021 - - - 35.8 31.3 37.6 44.1
+LP-BERT [27] 2022 1.18 100 - - - - -
+KGT5 [34] 2022 - - - 33.6 28.6 36.2 42.6
+CSPromp-KG [24] 2023 - - - 38 34.3 39.9 44.6
+ReSKGC [104] 2023 - - - 39.6 37.3 41.3 43.7
+Table 12: Experimental statistics of FB15k-237N dataset under the LP subtask in transductive setting.
+Model FB15k-237N
+MRR Hits@1 Hits@3 Hits@10
+TransE [52] 2013 25.5 15.2 30.1 45.9
+DistMult [71] 2014 20.9 14.3 23.4 33
+ComplEx [72] 2016 24.9 18 27.6 38
+ConvE [54] 2018 27.3 19.2 30.5 42.9
+RotatE [73] 2019 27.9 17.7 32 48.1
+CompGCN [93] 2020 31.6 23.1 34.9 48
+KG-BERT [19] 2019 20.3 13.9 20.1 40.3
+MTL-KGC [20] 2020 24.1 16 28.4 43
+GenKGC [32] 2022 - 18.7 27.3 33.7
+KG-S2S [35] 2022 35.4 28.5 38.8 49.3
+CSPromp-KG [24] 2023 36 28.1 39.5 51.1
+CD [38] 2024 37.2 28.8 41 53
+Table 13: Experimental statistics under the LP subtask in inductive setting on NELL-One dataset.
+Model Year NELL-One
+N-Shot MRR Hits@1 Hits@5 Hits@10
+GMatching [55] 2018 Five-Shot 20 14 26 31
+GMatching [55] 2018 One-Shot 19 12 26 31
+MetaR [105] 2019 Five-Shot 26 17 35 44
+MetaR [105] 2019 One-Shot 25 17 34 40
+StAR [25] 2021 Zero-Shot 26 17 35 45
+KG-S2S [35] 2022 Zero-Shot 31 22 41 48
+Table 14: Experimental statistics under the LP subtask in inductive setting on Wikidata5M dataset.
+Model Year Wikidata5M
+N-Shot MRR MR Hits@1 Hits@3 Hits@10
+DKRL [66] 2016 Zero-Shot 23.1 78 5.9 32 54.6
+RoBERTa [6] 2019 Zero-Shot 7.4 723 0.7 1 19.6
+KEPLER [28] 2021 Zero-Shot 40.2 28 22.2 51.4 73
+SimKGC [26] 2021 Zero-Shot 71.4 - 50.9 78.5 91.7
 Transformer-based models excel in capturing complex patterns and contextual information, significantly enhancing link prediction performance.
@@
 5 Future Directions
-The integration of large language models (LLMs) into knowledge representation learning (KRL) has demonstrated significant potential in enhancing the capabilities of artificial intelligence systems.
+The integration of LLM into KRL has demonstrated significant potential in enhancing the capabilities of artificial intelligence systems.
 However, there are numerous avenues for further research and development.
-This section outlines six promising directions to advance the field.
-5.1 Advanced Integration of LLMs and KRL for Enhanced Contextual Understanding
-One major area for future exploration is the advanced integration of LLMs with KRL to achieve deeper contextual understanding.
-Current models have shown that LLMs can significantly enhance the representation and reasoning capabilities of knowledge graphs (KGs).
+This section outlines five promising directions to advance the field.
+5.1 Advanced Integration of LLM and KRL for Enhanced Contextual Understanding
+One major area for future exploration is the advanced integration of LLM with KRL to achieve deeper contextual understanding.
+Current models have shown that LLM can significantly enhance the representation and reasoning capabilities of knowledge graphs.
 However, most existing approaches treat textual and structural data separately or integrate them at a superficial level.
@@
 For instance, developing models that can dynamically switch between textual and structural contexts based on the task requirements could lead to more robust and contextually aware systems.
-This might involve hybrid architectures that combine the strengths of LLMs for textual data with graph neural networks (GNNs) for structural data.
+This might involve hybrid architectures that combine the strengths of LLM for textual data with graph neural networks for structural data.
 These models could leverage attention mechanisms to weigh the importance of textual and structural information dynamically.
@@
 5.2 Development of Efficient Training and Inference Techniques
-The computational demands of LLMs are a significant challenge, particularly when integrating them with KRL tasks.
+The computational demands of LLM are a significant challenge, particularly when integrating them with KRL tasks.
 Future research should focus on developing more efficient training and inference techniques to make these models more accessible and practical for real-world applications.
-Techniques such as model pruning, quantization, and knowledge distillation have shown promise in reducing the computational load of LLMs without significantly sacrificing performance.
+Techniques such as model pruning, quantization, and knowledge distillation have shown promise in reducing the computational load of LLM without significantly sacrificing performance.
 Applying these techniques specifically to the context of KRL could yield models that are both powerful and efficient.
@@
 Instead of training models from scratch for each new task or domain, leveraging pre-trained models and fine-tuning them on specific KRL tasks can save significant computational resources.
-Continual learning techniques, where models incrementally learn from new data without forgetting previously acquired knowledge, can also enhance the adaptability and efficiency of LLM-based KRL systems.
+Continual learning techniques, where models incrementally learn from new data without forgetting previously acquired knowledge, can also enhance the adaptability and efficiency of LLM enhanced KRL systems.
 5.3 Improving Generalization and Robustness in KRL Models
-While LLM-enhanced KRL models have achieved impressive results on benchmark datasets, their generalization to unseen data and robustness in the face of noisy or incomplete data remain critical challenges.
+While LLM enhanced KRL models have achieved impressive results on benchmark datasets, their generalization to unseen data and robustness in the face of noisy or incomplete data remain critical challenges.
 Future research should aim to develop models that can generalize better across diverse domains and handle the inherent uncertainty and noise in real-world data.
@@
 5.4 Exploring Interpretability and Explainability in KRL Models
-As LLM-enhanced KRL models become more complex, their interpretability and explainability become increasingly important, especially in critical applications such as healthcare, finance, and legal domains.
+As LLM enhanced KRL models become more complex, their interpretability and explainability become increasingly important, especially in critical applications such as healthcare, finance, and legal domains.
 Future research should focus on developing methods to make these models more transparent and understandable.
-One approach is to develop techniques that can provide insights into the decision-making process of LLM-based KRL models.
+One approach is to develop techniques that can provide insights into the decision-making process of LLM enhanced KRL models.
 This could involve the use of attention mechanisms that highlight the most relevant parts of the input data, as well as the development of visualization tools that can illustrate how different parts of the model contribute to the final predictions.
@@
 5.5 Enhancing Human-AI Collaboration in KRL
-The integration of LLMs with KRL offers significant potential for enhancing human-AI collaboration.
+The integration of LLM with KRL offers significant potential for enhancing human-AI collaboration.
 Future research should explore how these advanced models can be leveraged to support and augment human decision-making and knowledge discovery processes.
-Developing interactive systems that allow humans to query and interact with LLM-enhanced KRL models can lead to more effective knowledge exploration and problem-solving.
+Developing interactive systems that allow humans to query and interact with LLM enhanced KRL models can lead to more effective knowledge exploration and problem-solving.
 For example, creating natural language interfaces that enable users to ask questions and receive explanations or suggestions from the model can make complex knowledge graphs more accessible and useful.
@@
 By fostering a symbiotic relationship between humans and AI, these systems can leverage the strengths of both to achieve superior performance.
-5.6 Addressing Ethical and Societal Implications
-As LLM-enhanced KRL models become more powerful and pervasive, it is crucial to address the ethical and societal implications of their deployment.
-Future research should focus on ensuring that these models are developed and used responsibly, with considerations for fairness, accountability, and transparency.
-One important aspect is to ensure that these models do not perpetuate or amplify biases present in the training data.
-Techniques such as bias detection and mitigation, as well as the development of fairness-aware algorithms, can help address these issues.
-It is also important to consider the broader societal impacts of deploying these models, including potential job displacement and the impact on privacy.
-Additionally, research should explore the development of regulatory frameworks and ethical guidelines for the use of LLM-enhanced KRL models.
-This could involve collaborations between researchers, policymakers, and industry stakeholders to ensure that these models are used in ways that align with societal values and ethical principles.
 6 Conclusion
-The integration of LLM with KRL represents a transformative advancement in AI, offering significant improvements in the ability to understand, represent, and utilize complex knowledge structures.
-This survey has systematically reviewed various approaches, categorizing them into encoder-based, encoder-decoder-based, and decoder-based methods, and evaluating their performance across multiple downstream tasks.
-Finally, there remain critical areas for further research.
+In this paper, we present a comprehensive survey of large language models (LLMs) that enhance Knowledge Representation Learning (KRL).
+We begin by delineating the utilization of LLMs in Knowledge Graphs (KGs) and summarize various Transformer architecture-based methodologies, including encoder-only, encoder-decoder, and decoder-only frameworks.
+These classifications facilitate an understanding of the evolution from purely structured representations to hybrid models that integrate structured and textual data.
+Additionally, we discuss the current research methodologies and their performance benefits in downstream tasks.
+We also outline available datasets, open-source codebases, and diverse applications.
+Conclusively, we propose several future research directions that are poised to drive advancements not only in KRL but also across broader fields within Natural Language Processing (NLP), leveraging the robust capabilities of LLMs.